### v1

# Quantum Universality in Composite Systems: A Trichotomy of Clifford Resources

Abstract
The efficient classical simulation of Clifford circuits constitutes a fundamental barrier to quantum advantage, typically overcome by injecting explicit non-Clifford “magic” resources. We demonstrate that for high-dimensional quantum systems (qudits), the resources required to break this barrier are strictly governed by the number-theoretic structure of the Hilbert space dimension $d$. By analyzing the adjoint action of the Clifford group, we establish a classification of single-qudit universality as a trichotomy. (I) For prime dimensions, the Clifford group is a maximal finite subgroup, and universality is robustly achieved by any non-Clifford gate. (II) For prime-power dimensions, the group structure fragments, requiring tailored diagonal non-Clifford gates to restore irreducibility. (III) Most notably, for composite dimensions with coprime factors, we demonstrate that standard entangling operations alone—specifically, generalized intra-qudit CNOT gates—generate the necessary non-Clifford resources to guarantee a dense subgroup of $\mathrm{SU}(d)$ without explicit diagonal magic injection. Our proofs rely on a new geometric criterion establishing that a subgroup with irreducible adjoint action is infinite if it contains a non-scalar element with projective distance strictly less than $1/2$ from the identity. These results suggest that “coprime architectures”—hybrid registers combining incommensurate dimensions—can bypass the standard overhead of magic state distillation.
I Introduction
What makes a quantum computer more powerful than a classical one? The standard narrative points to interference and entanglement, yet the Gottesman-Knill theorem reveals that highly-entangled states can be efficiently simulated classically when generated by Clifford operations alone [Gottesman1998, AaronsonGottesman2004]. The missing resource that lifts computation beyond classical reach is typically identified as “magic”: non-stabilizer states or non-Clifford gates that enable quantum universality—the ability to approximate any unitary operator to arbitrary precision by composing elements from a finite gate set [Kitaev1997, Deutsch1995]. This paper reveals that for quantum systems of composite dimension magic can emerge from the interplay of classical operations, with the local Hilbert space dimension unfolding as a previously unrecognized computational resource.
Consider a quantum register composed of subregisters with coprime local dimensions—say, a qutrit ($d=3$) coupled to a ququart ($d=4$). Coupling between them can be attained via the generalized intra-qudit controlled-addition gate: a permutation matrix that adds the control value to the target modulo $d$. Notably, this operation is classical and specified entirely by discrete modular arithmetic. Yet we prove that when such intra-qudit gates connect subregisters of coprime dimension, the group they generate together with local Clifford operations is dense in the special unitary group. Universality emerges from arithmetic alone.
This inverts the conventional understanding of quantum resources. In standard qubit architectures, the single-qubit Clifford group supplemented by CNOT gates remains finite—it is the multi-qubit Clifford group, efficiently simulable by the Gottesman-Knill theorem. Achieving universality demands a dedicated non-Clifford resource, such as the $T$-gate or magic states prepared through costly distillation protocols [Bravyi2005]. Our results show that this overhead is an artifact of homogeneous architectures built from identical subsystems, not an intrinsic feature of quantum computation. When local dimensions are coprime, the “free” entangling operations that any architecture provides are already sufficient.
The phenomenon has a transparent algebraic origin. The generalized CNOT is a classical reversible gate. However, when connecting subsystems of coprime dimensions $p$ and $q$, it acts on local Pauli operators to produce diagonal gates. By Bézout’s identity, these gates construct fractional phases $2\pi/pq$—inaccessible to either local Clifford group. Hence, the generation of dense subgroups follows from gates that individually preserve the computational basis. This suggests viewing local dimension as a computational resource. Unlike prime-power dimensions like $d=4$, coprime dimensions like $d=6$ allow the “mixing” of incompatible modular arithmetics, creating quantum resources from the arithmetic structure alone.
A finitely generated subgroup of $\mathrm{SU}(d)$ is dense if and only if it is infinite and acts irreducibly on the Lie algebra $\mathfrak{su}(d)$ via the adjoint representation. When applied to extensions of the Clifford group, this criterion identifies the specific obstructions to universality. In composite dimensions, unlike the prime case, the Clifford group is not maximal; therefore, a non-Clifford gate may generate a strictly larger yet still finite group. To rule out this possibility and ensure infiniteness, we introduce a geometric criterion: a subgroup with irreducible adjoint action is infinite if it contains a non-scalar element within projective distance $1/2$ of the identity.
Related work. The structure of universality for qubit systems is well understood: the Clifford group supplemented by any non-Clifford gate—typically the $T$-gate from the third level of the Clifford hierarchy—generates a dense subgroup of $\mathrm{SU}(2)$ [Gottesman1999, Nebe2001]. This “Clifford + TT” paradigm underlies modern fault-tolerant quantum computation, where magic states enable non-Clifford operations through gate teleportation [Bravyi2005, Knill2004]. For higher-dimensional systems, the situation is less complete. Universality criteria for qudits of prime dimension have been established [Howard2012, Campbell2012], and the maximality of the Clifford group in these cases follows from classical results on finite linear groups [Nebe2001]. However, composite dimensions with mixed prime factors have not been systematically studied, even as experimental platforms increasingly access non-binary levels [RingbauerEtAl2022, Low2025, Chicco_2023].
Universality for arbitrary dimension $d$, not based on the Clifford group, was established through both constructive and algebraic frameworks. Previous work proved universality by decomposing elements of $\mathrm{SU}(d^{n})$ into two-level unitaries, which were further expressed through phase and controlled operations [LuoWang2014]. Alternative constructions employed the generators of $\mathfrak{su}(d)$, where universal gate sets were generated via exponentiation of string operators, to produce $\mathrm{U}(d^{n})$ [VlasovJMP2002, VlasovSPIE2003].
Main results. We classify the resources required for single-qudit universality by the prime factorization of $d$:
  1. (I)
    Prime dimensions $\bm{(d=p).}$ The Clifford group acts irreducibly on the Lie algebra and is a maximal finite subgroup of $\mathrm{PSU}(d)$. We prove maximality holds if and only if $d$ is prime: adding any non-Clifford gate forces universality. In this regime, every non-Clifford gate is automatically magic.
  2. (II)
    Prime power dimensions $\bm{(d=p^{m},\,m\geq 2).}$ The adjoint representation becomes reducible. The Lie algebra fragments into invariant subspaces indexed by divisibility. Universality requires specific diagonal non-Clifford gates—members of the $T_{s}$ family whose phases restore irreducibility. Not every non-Clifford gate suffices; the arithmetic obstruction must be overcome by deliberate design.
Cases I and II share a common strategy: universality via diagonal non-Clifford gates (the $T_{s}$ family).
  1. (III)
    Composite dimensions with coprime factors $\bm{(d=d_{1}\cdots d_{n}).}$ Universality emerges from intra-qudit gates alone. When $d$ has pairwise coprime factors, generalized CNOTs connecting subsystems construct diagonal gates via Bézout’s identity. These gates satisfy both density conditions: they restore irreducibility by mixing the local Clifford structures, and they certify infiniteness by lying within projective distance $1/2$ of the identity. No explicit magic injection is required. In contrast to prime dimensions, the failure of maximality allows for intermediate finite groups, meaning that a gate can be non-Clifford without being magic.
Case III is our main contribution. In prime and prime-power dimensions (Case I and II), universality requires the explicit injection of diagonal non-Clifford gates ($T_{s}$). In contrast, coprime dimensions achieve universality using only generalized CNOTs. Although these are classical reversible permutations, they are non-Clifford resources. The arithmetic mismatch allows these permutation gates to generate a dense subgroup, making universality an intrinsic feature of the hardware that does not require explicit diagonal magic.
Structure of the paper. Section II establishes the algebraic framework: the density criterion, the structure of the generalized Pauli and Clifford groups, and the decomposition of composite-dimensional systems via the Chinese Remainder Theorem. Section III develops the geometric criterion for infiniteness and its spectral characterization. Sections IV, V, and VI analyze the three regimes of the trichotomy in detail. Section VII discusses implications of our findings and open problems.
II Preliminaries
II.1 Criteria for Universality
The practical realization of quantum algorithms relies on the ability to approximate arbitrary unitary transformations using a finite set of elementary gates. In the context of qudits, it is useful to distinguish between universality on a single local register and universality on a multi-qudit system. We adopt a topological definition based on the operator norm, which captures the physical requirement that any target unitary can be realized within an arbitrary error tolerance $\epsilon$.
Definition 1.
A finite set of gates $\mathcal{G}\subset\mathrm{SU}(d)$ is single-qudit universal if the group generated by $\mathcal{G}$, denoted $\langle\mathcal{G}\rangle$, is dense in $\mathrm{SU}(d)$ with respect to the operator norm topology. That is, for every $U\in\mathrm{SU}(d)$ and every $\epsilon>0$, there exists a finite sequence of gates $g_{1},\dots,g_{k}\in\mathcal{G}$ such that $\|U-g_{k}\cdots g_{1}\|<\epsilon$.
The focus of this work is primarily on establishing single-qudit universality. This focus is justified by the fact that the complexity of the universality problem resides almost entirely in the local structure. Once the single-qudit case has been completely characterized, extending this to universal quantum computation on an array of qudits requires only a single additional gate $V$ that breaks the preservation of the local tensor structure (an entangling resource). This reduction is formalized by the following seminal result:
Theorem 2 ([Brylinski2002]).
Let $\mathcal{G}_{\mathrm{local}}\subset\mathrm{SU}(d)$ be a single-qudit universal set and $V\in\mathrm{SU}(d^{2})$ be a two-qudit gate. Then, the set $\mathcal{G}_{\mathrm{local}}\cup\{V\}$ is universal for quantum computation on $n$ qudits if and only if $V$ cannot be decomposed into local operations, even up to a permutation of the qudits. Explicitly, universality is achieved if and only if for all $U_{1},U_{2}\in\mathrm{SU}(d)$:
$V\neq U_{1}\otimes U_{2}\quad\text{and}\quad V\neq(U_{1}\otimes U_{2})\cdot\mathrm{SWAP}.$ (1)
Consequently, establishing universality reduces to identifying when a finite set of gates generates a dense subgroup of the special unitary group. Proving density directly via topological arguments is often hard. The fundamental difficulty lies in distinguishing whether a finitely generated subgroup is dense or merely trapped within a proper, infinite closed Lie subgroup of lower dimension.
Instead, we employ an algebraic criterion based on the rigidity of simple Lie algebras. This criterion transforms the topological problem of density into two algebraic verifications: the irreducibility of the action on the associated Lie algebra and the infiniteness of the group.
To establish the notation, we introduce the necessary infinitesimal structures. The Lie algebra $\mathfrak{su}(d)$ consists of all trace-zero skew-Hermitian matrices:
[…]
$\operatorname{Ad}_{g}(X)=gXg^{\dagger},$ (3)
for $g\in\mathrm{SU}(d)$ and $X\in\mathfrak{su}(d)$. Since unitary conjugation preserves the Hilbert-Schmidt inner product, this action is unitary. A crucial property is that the irreducibility of this action is insensitive to complexification: the action on $\mathfrak{su}(d)$ is irreducible if and only if the complexified action on $\mathfrak{sl}(d,\mathbb{C})$ is irreducible.
We now establish the connection between the subgroup dimension and the representation.
Lemma 3.
Let $K$ be a closed subgroup of $\mathrm{SU}(d)$ with positive dimension. Then, the adjoint representation of $K$ on $\mathfrak{su}(d)$ is reducible unless $K=\mathrm{SU}(d)$.
Proof.
Let $\mathfrak{k}=\mathrm{Lie}(K)\subseteq\mathfrak{su}(d)$ be the Lie algebra of $K$. Since $K$ has positive dimension, $\mathfrak{k}\neq\{0\}$. The Lie algebra of a subgroup is invariant under the adjoint action of the group. Thus, $\mathfrak{k}$ is a $K$-invariant subspace of $\mathfrak{su}(d)$. If we assume the adjoint action of $K$ on $\mathfrak{su}(d)$ is irreducible, the only non-zero invariant subspace can be the entire space itself. Therefore, we must have $\mathfrak{k}=\mathfrak{su}(d)$.
Since $\mathrm{SU}(d)$ is connected and $K$ is a closed subgroup possessing the full Lie algebra of the ambient group, the connected component of the identity $K^{\circ}$ must coincide with $\mathrm{SU}(d)$. Since $K\subseteq\mathrm{SU}(d)$, it follows that $K=\mathrm{SU}(d)$. ∎
This lemma leads to one criteria for single-qudit universality.
Theorem 4.
Let $\mathcal{A}\subset\mathrm{SU}(d)$ be a finite set of gates. Then $\mathcal{A}$ is single-qudit universal if and only if:
  1. 1.
    The group $\langle\mathcal{A}\rangle$ is infinite, and
  2. 2.
    The adjoint representation of $\langle\mathcal{A}\rangle$ on $\mathfrak{su}(d)$ (equivalently on $\mathfrak{sl}(d,\mathbb{C})$) is irreducible.
Proof.
Let $K=\overline{\langle\mathcal{A}\rangle}$ be the topological closure of the generated group in $\mathrm{SU}(d)$.
$(\Rightarrow)$ Suppose $K=\mathrm{SU}(d)$. Since $\mathrm{SU}(d)$ has positive dimension, $\langle\mathcal{A}\rangle$ must be infinite. Furthermore, since $\mathfrak{su}(d)$ is a simple Lie algebra (for $d\geq 2$), the adjoint action of $\mathrm{SU}(d)$ is irreducible. Since $\langle\mathcal{A}\rangle$ is dense in $K$, any subspace invariant under $K$ is invariant under $\langle\mathcal{A}\rangle$. Thus, the dense subgroup must also act irreducibly.
$(\Leftarrow)$ Suppose $\langle\mathcal{A}\rangle$ is infinite and acts irreducibly on $\mathfrak{su}(d)$. Since the subgroup is infinite and contained in a compact group, its closure $K$ cannot be a finite set. Being a closed subgroup of a Lie group, $K$ is itself a Lie group with $\dim(K)>0$. The irreducibility of $\langle\mathcal{A}\rangle$ implies that the adjoint action of $K$ on $\mathfrak{su}(d)$ is also irreducible. We now apply Lemma 3 to $K$. Since $K$ has positive dimension and acts irreducibly, we conclude $K=\mathrm{SU}(d)$. ∎
In practice, we will often verify the second condition using the complex Lie algebra $\mathfrak{sl}(d,\mathbb{C})$, which simplifies calculations by allowing the use of complex bases such as the generalized Pauli operators.
II.2 The Clifford Group and its Structure
[…]
The generalized Pauli operators $X$ and $Z$ are defined by their action on the computational basis:
[…]
where arithmetic in the indices is taken modulo $d$, and $\omega=e^{2\pi i/d}$ is the primitive $d$-th root of unity. These operators satisfy the commutation relation
$ZX=\omega XZ.$
The Heisenberg group $H(d)$ (sometimes called the generalized Pauli group or just the Pauli group) consists of:
$H(d)=\begin{cases}\{\omega^{l}X^{a}Z^{b}:a,b,l\in\mathbb{Z}_{d}\}&\text{if }d\text{ is odd},\\ \{\zeta^{l}X^{a}Z^{b}:a,b\in\mathbb{Z}_{d},\,l\in\mathbb{Z}_{2d}\}&\text{if }d\text{ is even},\end{cases}$
where $\zeta=e^{\pi i/d}$ is a primitive $2d$-th root of unity. It is convenient to label the Pauli operators by vectors $u=(a,b)\in\mathbb{Z}_{d}\times\mathbb{Z}_{d}$, denoting $V_{u}=X^{a}Z^{b}$. The set $\{V_{u}:u\neq(0,0)\}$ forms an orthogonal basis for the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$ with respect to the Hilbert-Schmidt inner product.
Projective vs. Finite Clifford Groups
In the quantum information literature, the term “Clifford group” is often used interchangeably to refer to distinct mathematical objects. It is relevant for our analysis to distinguish between the projective group and its finite unitary representations.
Definition 5.
The (projective) single-qudit Clifford group, denoted $C(d)$, is defined as the normalizer of the Heisenberg group in the unitary group, modulo global phases:
$\begin{split}C(d)&=N_{\mathrm{U}(d)}(H(d))/\mathrm{U}(1)\\ &=\{U\in\mathrm{U}(d):UH(d)U^{\dagger}=H(d)\}/\mathrm{U}(1).\end{split}$ (4)
The adjoint action of any representative $U\in C(d)$ maps Pauli operators to Pauli operators, inducing an automorphism on the index space. This action corresponds to the special linear group over the ring $\mathbb{Z}/d\mathbb{Z}$. This structural rigidity is characterized by the short exact sequence [Tolar_2018]:
$1\longrightarrow\mathbb{Z}_{d}^{2}\longrightarrow C(d)\xrightarrow{\Psi}\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})\longrightarrow 1.$ (5)
Here, the kernel $\mathbb{Z}_{d}^{2}\cong H(d)/\mathrm{U}(1)$ corresponds, modulo global phases, to the Pauli operators themselves. Since elements of the Heisenberg group commute up to a scalar, their adjoint action on the Pauli basis is trivial on the index space (mapping $V_{u}\mapsto\omega^{k}V_{u}$ without changing $u$). In contrast, the quotient $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ characterizes the non-trivial structural changes: these gates actively transform the Pauli set by implementing group automorphisms on the indices, $u\mapsto\Psi(u)$.
While $C(d)$ is the natural object for studying symmetries, it is an abstract quotient group. The full normalizer $N_{\mathrm{U}(d)}(H(d))$ within the unitary group is continuous, as it includes the circle group $\mathrm{U}(1)$ of global phases. To properly formulate the universality problem—which asks whether a finite set of gates generates a dense subgroup—we must identify a concrete finite subgroup of unitary matrices that surjects onto $C(d)$.
Definition 6.
We define the finite single-qudit Clifford group, denoted $\mathcal{C}_{d}\subset\mathrm{U}(d)$, as the subgroup generated by the operator $X$, the generalized Hadamard gate $H_{d}$, and the phase gate $P$:
[…]
where $\rho_{d}=0$ if $d$ is even and $\rho_{d}=1$ if $d$ is odd.
To validate that the finite group $\mathcal{C}_{d}=\langle X,H_{d},P\rangle$ serves as the correct lifting of the projective Clifford group, we analyze the action of its generators under the homomorphism $\Psi$. The images of the Hadamard and phase gates correspond to the standard generators of the special linear group over the ring $\mathbb{Z}/d\mathbb{Z}$:
$\Psi(H_{d})=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\qquad\Psi(P)=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}$.
Since these matrices generate $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$, the group $\mathcal{C}_{d}$ surjects onto the symplectic quotient. Furthermore, the relation $Z=H_{d}XH_{d}^{\dagger}$ implies that $\mathcal{C}_{d}$ contains the full discrete Heisenberg group $H(d)$ (up to the phases required for group closure). Consequently, $\mathcal{C}_{d}$ fully captures the exact sequence structure defined in Eq. (5).
Remark 7.
Although we will often refer to $\mathcal{C}_{d}$ simply as “the Clifford group” by abuse of notation, it is distinct from the projective group $C(d)$. The group $\mathcal{C}_{d}$ fits into the short exact sequence:
$1\longrightarrow H(d)\longrightarrow\mathcal{C}_{d}\longrightarrow\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})\longrightarrow 1.$ (6)
Unlike the full normalizer in $\mathrm{U}(d)$, which contains the continuous circle group $\mathrm{U}(1)$, $\mathcal{C}_{d}$ is a finite group. Its order is strictly determined by the size of the $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ group and the scalar phases intrinsic to $H(d)$:
$|\mathcal{C}_{d}|=|H(d)|\cdot|\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})|.$ (7)
An important property for our analysis of composite systems is that this structure factorizes over coprime dimensions, allowing us to treat prime-power sectors independently.
Proposition 8.
Let $d=d_{1}\cdots d_{n}$, where the factors are pairwise coprime ($\gcd(d_{i},d_{j})=1$ for $i\neq j$). Then there is a group isomorphism:
$\mathcal{C}_{d}\cong\mathcal{C}_{d_{1}}\times\cdots\times\mathcal{C}_{d_{n}}.$
Proof.
It suffices to establish the case $n=2$ with coprime factors $d_{1},d_{2}$. We define the map $\Phi:\mathcal{C}_{d_{1}}\times\mathcal{C}_{d_{2}}\to\mathrm{U}(d_{1}d_{2})$ via the tensor product $\Phi(U_{1},U_{2})=U_{1}\otimes U_{2}$. We show this maps generators to generators, inducing an isomorphism onto $\mathcal{C}_{d}$.
Recall that by the Chinese Remainder Theorem (CRT), the map $\chi:\mathbb{Z}/d\mathbb{Z}\to\mathbb{Z}/{d_{1}}\mathbb{Z}\times\mathbb{Z}/{d_{2}}\mathbb{Z}$ given by $x\mapsto(x\bmod d_{1},x\bmod d_{2})$ is a ring isomorphism.
Since the Clifford generators are defined entirely by arithmetic operations on the basis indices, the CRT implies their factorization:
  1. 1.
[…]
  2. 2.
    Hadamard ($\bm{H}$): The Fourier transform kernel $\omega^{xy}$ factorizes because the inner product in $\mathbb{Z}_{d}$ decomposes into the sum of local inner products (modulo integer scaling), yielding $H_{d}\cong H_{d_{1}}\otimes H_{d_{2}}$.
  3. 3.
    Phase Gate ($\bm{P}$): The gate applies a phase $\omega^{Q(x)}$, where $Q(x)=x(x+\rho_{d})/2$ is a quadratic polynomial. Since $\chi$ is a ring homomorphism, it preserves polynomial evaluation: computing $Q(x)$ modulo $d$ is equivalent to computing pairs $(Q(x_{1})\bmod d_{1},Q(x_{2})\bmod d_{2})$. Consequently, the diagonal operator factorizes as $P_{d}\cong P_{d_{1}}\otimes P_{d_{2}}$.
Since the generators map bijectively and the groups have the same order, $\Phi$ is an isomorphism. ∎
III A Geometric Criterion for Infiniteness
Proving that a finitely generated matrix group is infinite is, in general, a non-trivial algebraic problem [detinko2013deciding]. While the irreducibility of the adjoint action ensures that the group is not trapped in a proper Lie subgroup of lower dimension, it does not rule out the possibility that the group is finite. To address this, we introduce a topological condition based on the discrete nature of finite groups.
Since we are interested in subgroups of $\mathrm{SU}(d)$, but physical operations are equivalent up to global phases, the appropriate topological setting is the projective unitary group $\text{PU}(d)=\mathrm{SU}(d)/Z(\mathrm{SU}(d))\cong\mathrm{U}(d)/\mathrm{U}(1)$. To capture the proximity to the identity in this quotient space, we employ a projective operator distance.
Definition 9.
The projective distance of $U\in\mathrm{SU}(d)$ is defined as:
$\mathrm{dist}_{\mathrm{proj}}(U,I)\coloneqq\min_{\phi\in[0,2\pi)}\|e^{i\phi}U-I\|,$ (8)
[…]
Note that the minimization is well-defined because the function $f(\phi)=\|e^{i\phi}U-I\|$ is continuous on the compact interval $0,2\pi]$. Also, while $U\in\mathrm{SU}(d)$, the optimal representative $e^{i\phi}U$ may lie in $\mathrm{U}(d)$.
Theorem 10.
Let $G\subseteq\mathrm{SU}(d)$ be a subgroup that acts irreducibly on the Lie algebra $\mathfrak{su}(d)$ via the adjoint representation. Then, $G$ is infinite if and only if there exists an element $h\in G$ such that:
$0<\mathrm{dist}_{\mathrm{proj}}(h,I)<\frac{1}{2}.$ (9)
The lower bound (>0>0) guarantees that $h$ is not a scalar, while the upper bound (<1/2<1/2) places it within a neighborhood of the identity prohibited for finite primitive groups.
III.1 Proof of the Criterion
To prove Theorem 10, we first establish a lemma connecting the irreducibility of the adjoint representation to the structure of normal abelian subgroups. Recall that a linear group $G\subseteq\mathrm{GL}(V)$ is called imprimitive if it preserves a decomposition of the space $V=\bigoplus V_{i}$ into subspaces permuted by $G$. If no such decomposition exists and $G$ is irreducible, it is called primitive.
Lemma 11.
Let $G\subseteq\mathrm{SU}(d)$ be a subgroup. If the adjoint representation of $G$ on $\mathfrak{su}(d)$ is irreducible, then $G$ acts primitively on $\mathbb{C}^{d}$, and every normal abelian subgroup of $G$ is contained in the center $Z(\mathrm{SU}(d))=\{\omega I:\omega^{d}=1\}$.
Proof.
First, we show that irreducibility of the adjoint action implies primitivity on $\mathbb{C}^{d}$. Suppose $G$ is not primitive. Then it acts either reducibly or imprimitively on $\mathbb{C}^{d}$:
  • If $G$ acts reducibly, there exists a proper invariant subspace $W\subset\mathbb{C}^{d}$. The subalgebra $\mathfrak{k}=\{X\in\mathfrak{su}(d):XW\subseteq W\}$ would then be a proper non-zero invariant subspace for the adjoint representation.
  • If $G$ is imprimitive, preserving a decomposition $V=\bigoplus_{i=1}^{m}V_{i}$, then the subalgebra $\mathfrak{h}=\{X\in\mathfrak{su}(d):XV_{i}\subseteq V_{i}\,\forall i\}$ is a non-zero subspace invariant under conjugation by $G$.
In both cases, the adjoint representation would be reducible, contradicting the hypothesis. Thus, $G$ is primitive.
Now, let $A\trianglelefteq G$ be a normal abelian subgroup. Since elements of $A$ commute, they share a simultaneous eigenspace decomposition $V=\bigoplus_{\chi}V_{\chi}$. Since $A$ is normal, $G$ permutes these eigenspaces. If there were more than one distinct eigenspace, this would constitute a system of imprimitivity. Since $G$ is primitive, there must be only one component $V=V_{\chi}$. This implies that every $a\in A$ acts as a scalar $a=\chi(a)I$, so $A\subseteq Z(SU(d))$. ∎
We can now prove the main theorem.
Proof of Theorem 10.
$(\Rightarrow)$ Suppose $G$ is infinite. Since the adjoint representation is irreducible, Theorem 4 implies that $G$ is dense in $\mathrm{SU}(d)$ with respect to the norm topology.
We define the function $F:\mathrm{SU}(d)\to\mathbb{R}^{\geq 0}$ by $F(U)=\mathrm{dist}_{\mathrm{proj}}(U,I)$. This function is continuous, as it is the composition of continuous operations. Consider the set:
$\mathcal{U}=\{U\in\mathrm{SU}(d):F(U)<1/2\}=F^{-1}\big([0,1/2)\big).$
Since $F$ is continuous, $\mathcal{U}$ is an open set in $\mathrm{SU}(d)$. Furthermore, $\mathcal{U}$ is non-empty as it contains the identity ($F(I)=0$).
A dense subgroup $G$ must intersect any non-empty open set infinitely many times. Therefore, $G\cap\mathcal{U}$ contains infinitely many elements. We can thus choose an element $h\in G\cap\mathcal{U}$ that is not a scalar multiple of the identity (i.e., $F(h)>0$). This $h$ satisfies the condition $0<\mathrm{dist}_{\mathrm{proj}}(h,I)<1/2$.
$(\Leftarrow)$ We proceed by contradiction. Suppose $G$ is finite and contains $h$ with $0<\mathrm{dist}_{\mathrm{proj}}(h,I)<1/2$. Let $\phi$ be the phase achieving the minimum distance, and define $h^{\prime}=e^{i\phi}h$.
Consider the extended group $G^{\prime}=\langle G,e^{i\phi}I\rangle$. We claim that $G^{\prime}$ is finite. Since $G$ is finite, $h$ has finite order, implying its eigenvalues are roots of unity. The minimal arc containing the spectrum is delimited by two such eigenvalues. Consequently, the centering phase $\phi$, which corresponds to the midpoint of this arc, is necessarily a rational multiple of $2\pi$. Thus, the scalar generator $e^{i\phi}I$ has finite order. Since $G$ is finite and $e^{i\phi}I$ is a central element of finite order, the generated group $G^{\prime}$ is finite.
We now rely on a structural property established in the proof of Theorem 5.7 in Ref. Dixon1971: for any finite subgroup $G^{\prime}\subset\mathrm{GL}(d,\mathbb{C})$, the subgroup $A=\langle\{g\in G^{\prime}:\|g-I\|<1/2\}\rangle$ is a normal abelian subgroup. Since $h^{\prime}\in G^{\prime}$ satisfies $\|h^{\prime}-I\|<1/2$, it implies $h^{\prime}\in A$, so $A$ is non-trivial. However, since the adjoint representation of $G^{\prime}$ (which coincides with that of $G$) is irreducible, Lemma 11 forces any normal abelian subgroup to be central. Thus, $h^{\prime}$ must be a scalar matrix. This implies $h$ is a scalar, which contradicts the assumption that $\mathrm{dist}_{\mathrm{proj}}(h,I)>0$. Therefore, $G$ cannot be finite. ∎
III.2 Spectral Characterization via Projective Distance
In practice, we verify the criterion by inspecting the eigenvalues. To adapt our infiniteness criterion to the projective setting, we identify the group of physical phases $U(1)$ with the quotient group $\mathbb{R}/\mathbb{Z}$ via the exponential map $\theta\mapsto e^{2\pi i\theta}$.
To evaluate the projective distance efficiently, we first introduce a geometric measure of the angular concentration of the eigenvalues.
Definition 12.
Let $h\in\mathrm{U}(d)$ be a unitary matrix. We define the spectral span $\ell(h)$ as the minimum $\Delta\in[0,1]$ such that there exists $\phi_{0}\in\mathbb{R}$ and $\delta_{j}\in[-\Delta/2,\Delta/2]$ satisfying:
$\sigma(h)=\left\{e^{2\pi i(\phi_{0}+\delta_{j})}\right\}_{j=1}^{d}.$ (10)
This quantity $\ell(h)$ represents the “angular width” of the spectrum on the unit circle. The following lemma establishes that the projective distance is determined uniquely by $\ell(h)$.
Lemma 13.
For any unitary matrix $h\in\mathrm{U}(d)$, the projective distance to the identity is given by:
$\mathrm{dist}_{\mathrm{proj}}(h,I)=2\sin\left(\frac{\pi\ell(h)}{2}\right).$ (11)
Proof.
Recall that $\mathrm{dist}_{\mathrm{proj}}(h,I)=\min_{\phi}\max_{j}|e^{2\pi i(\theta_{j}-\phi)}-1|$. Using the chordal formula $|e^{i\alpha}-1|=2|\sin(\alpha/2)|$, this is equivalent to minimizing the maximum angular deviation of the rotated eigenvalues from the identity (0).
Let $\ell(h)$ be the spectral span as per Definition 12. By existence of the minimum, we can choose the optimal parameters $\phi_{0}$ and $\{\delta_{j}\}$ such that the eigenvalues are $\lambda_{j}=e^{2\pi i(\phi_{0}+\delta_{j})}$ with maximal deviation $\max|\delta_{j}|=\ell(h)/2$. We choose the global phase shift $\phi=\phi_{0}$ (rotation by $\phi_{0}$). This centers the spectrum around 0, so the angular phases become exactly $2\pi\delta_{j}$. The operator norm is determined by the eigenvalue farthest from $1$:
$\|e^{-2\pi i\phi_{0}}h-I\|=\max_{j}2\left|\sin\left(\frac{2\pi\delta_{j}}{2}\right)\right|=2\sin\left(\pi\frac{\ell(h)}{2}\right).$
Any other interval covering the spectrum has length $\ell^{\prime}\geq\ell(h)$, implying a maximum deviation of at least $\ell^{\prime}/2$, which would yield a larger distance. Thus, the minimum is attained at the spectral span. ∎
Corollary 14.
A subgroup $G\subseteq\mathrm{SU}(d)$ with irreducible adjoint action is infinite if it contains a gate $h$ with spectral span $\ell(h)$ satisfying:
$0<\pi\ell(h)<2\arcsin(1/4)\approx 0.5053.$ (12)
IV Case I: Prime Dimensions
We begin our analysis with the simplest case: dimensions $d=p$ where $p$ is a prime number. In this setting, the Clifford group exhibits a remarkable structural rigidity. To understand this, we must first establish a general correspondence between the group’s action on the Lie algebra and the symplectic geometry of the discrete phase space. This framework will prove crucial not only for establishing universality in prime dimensions but also for diagnosing the obstructions that arise in prime-power dimensions (Sec. V).
IV.1 The Adjoint Action and $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ Orbits
To explicitly decompose the adjoint representation of the Clifford group $C(d)$ on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$ into a direct sum of irreducible subrepresentations, we employ the theory of representations of group extensions developed by Clifford [clifford1937, Serre1977].
The definition of the Clifford group implies that its adjoint action is monomial with respect to the generalized Pauli basis: elements of the group permute the basis vectors $\{V_{u}\}$ while applying phase factors. This structure naturally invites the application of Clifford’s theory to the short exact sequence characterizing the group (Eq. 5).
The Pauli operators $\{V_{u}:u\in\mathbb{Z}_{d}^{2}\setminus\{0\}\}$ form an orthogonal basis for the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$ with respect to the Hilbert-Schmidt inner product. We analyze the adjoint representation $\operatorname{Ad}:C(d)\to\mathrm{GL}(\mathfrak{sl}(d,\mathbb{C}))$. According to Clifford’s theory, the restriction of the representation to the normal subgroup $\mathbb{Z}_{d}^{2}$ decomposes the space into a direct sum of isotypic components associated with the orbits of the quotient group acting on the character group $\widehat{\mathbb{Z}_{d}^{2}}$.
In our case, the Pauli basis elements $V_{u}$ form a simultaneous eigenbasis for the action of $\mathbb{Z}_{d}^{2}$. Explicitly, for any element $u\in\mathbb{Z}_{d}^{2}$, which corresponds to a class of Pauli operators $V_{v}]$, the adjoint action is given by the commutation relation:
$\operatorname{Ad}_{V_{v}}(V_{u})=V_{v}V_{u}V_{v}^{\dagger}=\omega^{\langle v,u\rangle}V_{u}.$
Thus, each 1-dimensional subspace spanned by $V_{u}$ corresponds to a distinct character $\chi_{u}\in\widehat{\mathbb{Z}_{d}^{2}}$ defined by $\chi_{u}(v)=\omega^{\langle v,u\rangle}$. The action of the quotient group $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ permutes these characters (and thus the basis vectors) according to the linear map on the indices $u\mapsto\Psi_{g}(u)$.
Consequently, the adjoint representation decomposes into a direct sum of irreducible $C(d)$-invariant subspaces determined by the $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ orbits:
$\mathfrak{sl}(d,\mathbb{C})=\bigoplus_{\mathcal{O}\in\mathbb{Z}_{d}^{2}/\mathrm{SL}(2,\mathbb{Z}_{d})}W_{\mathcal{O}},\quad\text{where}\quad W_{\mathcal{O}}=\bigoplus_{u\in\mathcal{O}}\mathbb{C}V_{u}.$
This leads to the following correspondence:
Lemma 15.
Let $d$ be any integer. The decomposition of the adjoint representation of $C(d)$ on $\mathfrak{sl}(d,\mathbb{C})$ into irreducible subspaces is in one-to-one correspondence with the orbits of the natural action of the group $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ on the non-zero index vectors $\mathbb{Z}_{d}^{2}\setminus\{0\}$. Specifically, for each orbit $\mathcal{O}$, the subspace $W_{\mathcal{O}}$ is an irreducible $C(d)$-module.
To apply this correspondence, we require an explicit classification of the $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ orbits. For dimensions of the form $d=p^{m}$ (which includes the prime case $m=1$), the orbits are fully determined by the divisibility of the index vector.
Lemma 16.
Let $d=p^{m}$ be a prime power. Under the action of $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ on the $\mathbb{Z}_{d}^{2}$, the quantity $g(u)=\gcd(u_{1},u_{2},d)$ is an invariant. Two vectors $u,v\in\mathbb{Z}_{d}^{2}$ belong to the same orbit if and only if they share this invariant:
$\mathcal{O}_{u}=\mathcal{O}_{v}\iff\gcd(u_{1},u_{2},d)=\gcd(v_{1},v_{2},d).$
Proof.
The greatest common divisor is preserved under integer linear combinations with unit determinant. If $v=Mu$ for some $M\in\mathrm{SL}(2,\mathbb{Z}_{d})$, then the components of $v$ are linear combinations of the components of $u$. Thus, any common divisor of $u_{1},u_{2}$ (and $d$) must also divide $v_{1},v_{2}$. Since $M$ is invertible, the reverse implication holds, ensuring $g(u)=g(v)$.
Now, we show that any vector $u$ with invariant $g(u)=p^{j}$ (for $0\leq j\leq m$) lies in the orbit of the canonical vector $e_{j}=(p^{j},0)^{T}$.
Write $u=(p^{j}a,p^{j}b)^{T}$. Since the total gcd with the modulus $d=p^{m}$ is $p^{j}$, the remaining factors must satisfy $\gcd(a,b,p)=1$. By Bézout’s identity, there exist integers $x,y$ such that $ax+by\equiv 1\pmod{p^{m}}$. We construct the matrix:
$M=\begin{pmatrix}x&y\\ -b&a\end{pmatrix}$.
Observe that $\det(M)=ax+by\equiv 1\pmod{p^{m}}$, so $M\in\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$. Applying this transformation to $u$:
$M\begin{pmatrix}p^{j}a\\ p^{j}b\end{pmatrix}=p^{j}\begin{pmatrix}x&y\\ -b&a\end{pmatrix}\begin{pmatrix}a\\ b\end{pmatrix}=p^{j}\begin{pmatrix}ax+by\\ -ba+ab\end{pmatrix}=\begin{pmatrix}p^{j}\\ 0\end{pmatrix}$.
Since any vector with invariant $p^{j}$ can be mapped to the same canonical state $(p^{j},0)^{T}$, all such vectors belong to the same orbit. ∎
IV.2 Irreducibility of the Adjoint Action
The first condition of Theorem 4 is the irreducibility of the action on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$. We can now prove this characterization using the tools developed.
Theorem 17.
The adjoint representation of the Clifford group $C(d)$ on $\mathfrak{sl}(d,\mathbb{C})$ is irreducible if and only if $d$ is prime.
Proof.
If $d=p$ is prime, the only divisors of $d$ are $1$ and $p$. For any non-zero vector $u\in\mathbb{Z}_{p}^{2}\setminus\{0\}$, the greatest common divisor is $\gcd(u,p)=1$. By Lemma 16, all such vectors belong to a single $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ orbit $\mathcal{O}_{1}$. Consequently, by Lemma 15, the representation space $\mathfrak{sl}(p,\mathbb{C})$ corresponds to a single irreducible module $W_{\mathcal{O}_{1}}$. Thus, the action is irreducible.
Conversely, if $d$ is composite, there exists a proper divisor $k$ ($1<k<d$). The set of vectors with $\gcd(u,d)=k$ forms an orbit disjoint from the set of vectors with $\gcd(u,d)=1$. This implies the existence of at least a proper non zero invariant subspaces, rendering the representation reducible. ∎
IV.3 Maximality and Universality
To establish universality for prime dimensions $d=p$, we rely on the fact that any extension of the Clifford group by a non-Clifford gate forces the group to become infinite. This property stems from the maximality of the Clifford group within the finite subgroups of the projective unitary group. While this fact is widely recognized in quantum information [Gottesman1999], a rigorous mathematical proof relies on deep results from finite group theory.
For the general case of systems with $n$-qudits ($d$ prime), Nebe, Rains, and Sloane [Nebe2001] established maximality using sophisticated number-theoretic techniques. Specifically, they utilized $l$-adic integrality and reduction modulo suitable rational primes to apply the classification of subgroups of finite general linear groups. However, as noted in Ref. [Nebe2001], for the single-qudit case (d=pd=p)—which is central to this work—the sufficiency of the primality condition follows from earlier work by Lindsey [Lindsey1970]. Unlike the general multi-qudit case, Lindsey’s proof relies primarily on character theory and the classification of linear groups of prime degree. Complementing this classical result, we demonstrate that primality is also a necessary condition, thereby establishing the following full characterization. For completeness, we provide a self-contained proof of the forward direction based on a Corollary in [Lindsey1970].
Theorem 18.
The single-qudit Clifford group $\mathcal{C}_{d}$ is a maximal finite subgroup of the unitary group modulo scalars (i.e., any finite group $G\supsetneq\mathcal{C}_{d}$ satisfies $G=\langle\mathcal{C}_{d},\lambda I\rangle$) if and only if $d$ is a prime number.
Proof.
The proof addresses the sufficient and necessary conditions separately.
1. Sufficiency. Let $p>5$ be prime. We will show that $\mathcal{C}_{p}$ is a maximal finite subgroup of $\mathrm{SU}(p)/U(1)$.
Recall the exact sequence
$1\longrightarrow H(d)\longrightarrow\mathcal{C}_{d}\longrightarrow\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})\longrightarrow 1,$
where $H(p)$ is the Heisenberg group (an extraspecial $p$-group of order $p^{3}$ and exponent $p$). The group $H(p)$ acts irreducibly on $\mathbb{C}^{p}$ and is a normal subgroup of $\mathcal{C}_{p}$.
Suppose, by contradiction, that there exists a finite group $G$ with $\mathcal{C}(p)\subsetneq G\subseteq\mathrm{U}(p)/\mathrm{U}(1)$, where the inclusion is strict and then $G$ is not obtained from $\mathcal{C}_{p}$ by merely adjoining scalar matrices.
The natural $p$-dimensional representation of $G$ is faithful, primitive, and irreducible (primitivity follows since $p$ prime and then the adjoint representation is irreducible).
We apply the following result of Lindsey:
Theorem 19 ([Lindsey1970, Corollary, p. 58]).
Let $G$ be a finite group with a faithful, primitive, irreducible representation of prime degree $p>5$. If $p^{4}\mid|G|$, then $G$ has a normal nonabelian $p$-subgroup $D$ of order $p^{3}$, and $G/D$ is isomorphic to a subgroup of $\mathrm{SL}(2,\mathbb{Z}/p\mathbb{Z})$.
Since $\mathcal{C}_{p}\subset G$ and $p^{4}\mid|\mathcal{C}_{p}|$, we have $p^{4}\mid|G|$. By Theorem 19, $G$ contains a normal nonabelian $p$-subgroup $D$ of order $p^{3}$ with $G/D\hookrightarrow\mathrm{SL}(2,\mathbb{Z}/p\mathbb{Z})$.
The group $D$ must coincide with $H(p)$: both are extraspecial $p$-groups of order $p^{3}$ acting irreducibly on $\mathbb{C}^{p}$, and such a group is unique up to conjugacy in $\mathrm{GL}(p,\mathbb{C})$.
Therefore, $G$ normalizes $H(p)$, which means $G$ is contained in the normalizer $N_{\mathrm{U}(p)}(H(p))$. But this normalizer, modulo scalars, is precisely the Clifford group $\mathcal{C}_{p}$. Hence $G=\mathcal{C}(p)$, contradicting our assumption.
For the remaining small primes $p\in\{2,3,5\}$, maximality follows from the explicit classification of finite primitive linear groups of degree $p$. If a finite extension $G\supsetneq\mathcal{C}_{p}$ existed, it would necessarily possess a Sylow $p$-subgroup of order at least $|\mathcal{C}_{p}|_{p}\geq p^{4}$ (accounting for the Heisenberg group and $SL(2,\mathbb{Z}/p\mathbb{Z})$ part). However, the classifications by Blichfeldt [Blichfeldt1917] and Brauer [Brauer1967] reveal that no primitive group—other than the normalizers of extraspecial groups—supports such a structure. For instance, for $p=5$, Brauer established that the only other primitive candidates involve central extensions of $A_{5},A_{6},$ or $\text{PSL}(2,11)$, whose Sylow $5$-subgroups have orders at most $25$, strictly less than the required order. Similar order constraints on the Sylow $p$-subgroups rule out extensions for $p=2$ and $p=3$.
2. Necessity ($d$ composite). We prove by contradiction: if $d$ is composite, $\mathcal{C}_{d}$ is not maximal. We distinguish two cases based on the prime factorization of $d$.
Case A: Prime Powers ($d=p^{m}$ with $m\geq 2$). We explicitly construct a strictly larger finite group by identifying a system of imprimitivity. Recall that the Pauli operators $X,Z$ generate the Heisenberg group. Consider the subgroup generated by the $p^{m-1}$-th powers of these generators:
$\mathcal{A}=\langle X^{p^{m-1}},Z^{p^{m-1}}\rangle.$
We verify the commutativity of these elements. The group commutator is determined by the phase $\omega=e^{2\pi i/p^{m}}$. Since $X^{p^{m-1}},Z^{p^{m-1}}]=\omega^{p^{2m-2}}I$, and $2m-2\geq m$ for $m\geq 2$, the commutator is the identity. Thus, $\mathcal{A}$ is a normal abelian subgroup.
The Hilbert space decomposes into the direct sum of the simultaneous eigenspaces of $\mathcal{A}$:
$\mathbb{C}^{d}=\bigoplus_{\lambda\in\widehat{\mathcal{A}}}V_{\lambda}.$
The index set $\Lambda=\widehat{\mathcal{A}}$ is isomorphic to $\mathbb{Z}_{p}^{2}$ and has cardinality $K=p^{2}$. The action of $\mathcal{C}_{d}$ by conjugation preserves $\mathcal{A}$, and therefore induces a permutation action on the set of indices $\Lambda$. This proves that $\mathcal{C}_{d}$ acts imprimitively.
To prove non-maximality, we analyze the embedding of $\mathcal{C}_{d}$ into the stabilizer of this decomposition. Let $\Phi:\mathcal{C}_{d}\to S_{\Lambda}$ be the homomorphism describing the induced permutation of the $K$ subspaces. We distinguish two subcases:
  1. 1.
    Failure by Permutation ($p\geq 3$): The image $\text{Im}(\Phi)$ corresponds to the affine symplectic group $\mathrm{ASL}(2,\mathbb{Z}_{p})$, which has order $p^{3}(p^{2}-1)$. For $p\geq 3$, this order is strictly smaller than that of the symmetric group $|S_{p^{2}}|=(p^{2})!$. Thus, there exists a permutation $\pi\in S_{\Lambda}\setminus\text{Im}(\Phi)$. We can strictly extend $\mathcal{C}_{d}$ by adding a finite-order unitary $U_{\pi}$ that permutes the subspaces according to $\pi$ (acting as identity within each block). Note that $U_{\pi}$ is not a scalar multiple of the identity, as it performs a non-trivial permutation of orthogonal subspaces.
  2. 2.
    Failure by Phase Independence ($p=2$): In this case, $K=4$ and the permutation groups coincide ($|\mathrm{ASL}(2,2)|=|S_{4}|=24$). However, maximality fails in the kernel of $\Phi$ (the subgroup fixing the blocks). Restricted to the action on the set of blocks, the diagonal elements of the Clifford group induce relative phases of the form $\phi(u)=\omega^{Q(u)}$ for $u\in\Lambda$, where $Q$ is necessarily a polynomial of degree at most 2 (a quadratic form). In contrast, the stabilizer of the decomposition allows for arbitrary phase assignments $\phi(u)=\omega^{f(u)}$. Since the space of all functions on $\mathbb{Z}_{2}^{2}$ is strictly larger than the space of quadratic forms, there exists a diagonal unitary $D$ (e.g., $D=\mathrm{diag}(1,1,1,i)$ relative to the blocks) which belongs to the finite stabilizer but not to $\mathcal{C}_{d}$. This operator is clearly not a scalar multiple of the identity, as it possesses distinct eigenvalues.
In both instances, we construct a finite group $G^{\prime}=\langle\mathcal{C}_{d},U_{\text{ext}}\rangle$ such that $\mathcal{C}_{d}\subsetneq G^{\prime}$ and $G^{\prime}\neq\langle\mathcal{C}_{d},Z(G^{\prime})\rangle$. Thus, $\mathcal{C}_{d}$ is not maximal.
Case B: Coprime Factors ($d=a\cdot b$ with $\gcd(a,b)=1$). In this case, the failure of maximality arises from the tensor product structure. By Proposition 8, the group decomposes as $\mathcal{C}_{d}\cong\mathcal{C}_{a}\times\mathcal{C}_{b}$. This group acts on $\mathcal{H}\cong\mathbb{C}^{a}\otimes\mathbb{C}^{b}$ and preserves the tensor decomposition. We can strictly extend this group by adding finite-order unitary operators that normalize the tensor structure but do not belong to the direct product. For example, if $a=b$, the SWAP operator $S$ has finite order and normalizes the group, but $S\notin\mathcal{C}_{d}$. Even if $a\neq b$, the group is contained in the finite normalizer of the product structure, which allows for symmetries larger than the restricted $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ action induced by $\mathcal{C}_{d}$. ∎
IV.4 Universality and Diagonal Resources
The combination of irreducibility (Theorem 17) and maximality (Theorem 18) yields an immediate criterion for universality, valid only in prime dimensions.
Corollary 20.
Let $p$ be a prime. Let $g\in\mathrm{SU}(p)$ be any unitary gate. If $g\notin\mathcal{C}_{p}$ (modulo scalars), then the set $\mathcal{C}_{p}\cup\{g\}$ is single-qudit universal.
This corollary shifts the burden of finding universal sets to simply identifying a single non-Clifford gate. The most natural candidates are diagonal gates.
Definition 21.
Let $\zeta:\mathbb{Z}_{d}\to\mathrm{U}(1)$ be a phase function. We define the associated diagonal gate $T_{\zeta}$ by its action on the computational basis:
$T_{\zeta}|x\rangle=\zeta(x)|x\rangle,\quad\text{for }x\in\mathbb{Z}_{d}.$ (13)
To determine whether a specific $T_{\zeta}$ belongs to the Clifford group, we examine its conjugation action on the operator $X$. A simple calculation shows that $T_{\zeta}XT_{\zeta}^{\dagger}=DX$, where $D$ is a diagonal operator encoding the relative phases. This motivates the introduction of the differential of the phase function.
Definition 22.
The differential (or 1-coboundary) of a function $\zeta:\mathbb{Z}_{d}\to\mathrm{U}(1)$ is the map $\delta_{\zeta}:\mathbb{Z}_{d}\times\mathbb{Z}_{d}\to\mathrm{U}(1)$ defined by:
$\delta_{\zeta}(x,y)=\frac{\zeta(x+y)}{\zeta(x)\zeta(y)}.$ (14)
Recall that a function $b:\mathbb{Z}_{d}\times\mathbb{Z}_{d}\to\mathrm{U}(1)$ is called a bicharacter if it is linear in both arguments.
Lemma 23.
Let $d$ be any integer. A diagonal gate $T_{\zeta}$ belongs to the Clifford group $\mathcal{C}_{d}$ if and only if its differential $\delta_{\zeta}$ is a bicharacter.
Proof.
The conjugation action on the shift operator $X^{y}$ is given by:
$T_{\zeta}X^{y}T_{\zeta}^{\dagger}|k\rangle=\frac{\zeta(k)}{\zeta(k-y)}X^{y}|k\rangle.$
For the result to be a generalized Pauli operator (of the form $Z^{a}X^{y}$), the coefficient function must be a character (linear in $k$). In terms of the differential, this condition requires $\delta_{\zeta}(k,-y)$ to be linear in $k$. Since this must hold for all $y$, $\delta_{\zeta}$ must be bilinear.
Note that for the cyclic group $\mathbb{Z}_{d}$, any symmetric bicharacter takes the simple form $b(x,y)=\omega^{cxy}$, where $\omega=e^{2\pi i/d}$ is the primitive $d$-th root of unity and $c$ is an integer constant. Thus, the condition implies that $\zeta(x)$ must be a quadratic phase polynomial $\omega^{ax^{2}+bx}$. ∎
This lemma provides a constructive method to generate universal gate sets. For a prime dimension $p$, any diagonal gate whose differential is not a bicharacter (corresponding to a non-quadratic phase) is universal.
IV.5 The $T_{s}$-gate Family
The criteria for universality derived in the previous sections—maximality for prime dimensions and irreducibility via mixing for composite dimensions—share a common requirement: the availability of a unitary that lies strictly outside the local Clifford group. To characterize such operators, for any dimension $d$, we introduce a generic family of diagonal gates.
Definition 24 (The $T_{s}$-gate Family).
For any dimension $d\geq 2$ and any integer resolution $s\geq 2$, we define the diagonal gate $T_{s}$ acting on the computational basis as:
$T_{s}=\sum_{k\in\mathbb{Z}_{d}}\exp\left(\frac{2\pi ik}{s}\right)|k\rangle\langle k|.$ (15)
This family generalizes the standard Clifford hierarchy. For instance, $T_{d}$ recovers the Pauli $Z$ gate (order $d$), and $T_{2d}$ corresponds to the generalized phase gate.
To determine the “Cliffordness” of $T_{s}$ for an arbitrary dimension, we analyze its conjugation action on the Pauli shift operator $X$. A direct calculation yields:
$\operatorname{Ad}_{T_{s}}(X)=T_{s}XT_{s}^{\dagger}=\exp\left(\frac{2\pi i}{s}\right)X.$ (16)
For $T_{s}$ to belong to the Clifford group $\mathcal{C}_{d}$, it must normalize the Heisenberg group. Since the result is proportional to the operator $X$, this condition holds if and only if the $e^{2\pi i/s}I$ is in the Heisenberg group:
  • Odd Dimensions: The phases are generated by the primitive $d$-th root of unity $\omega=e^{2\pi i/d}$. Thus, $T_{s}\in\mathcal{C}_{d}\iff s\mid d$.
  • Even Dimensions: The Clifford group generates phases of order $2d$. Specifically for qubits ($d=2$), the group includes the order-4 phase $i$. Thus, generally, $T_{s}\in\mathcal{C}_{d}\iff s\mid 2d$.
The previous discussion leads to a unified criterion for when $T_{s}$ constitutes a non-Clifford resource.
Proposition 25.
The gate $T_{s}$ is a non-Clifford resource for $\mathrm{SU}(d)$ if and only if the order $s$ satisfies:
$s\nmid K_{d},\quad\text{where }K_{d}=\begin{cases}d&\text{if }d\text{ is odd},\\ 2d&\text{if }d\text{ is even}.\end{cases}$ (17)
Corollary 26.
Let $p$ be a prime number. The set $\mathcal{C}_{p}\cup\{T_{s}\}$ is single-qudit universal if and only if $s\nmid K_{p}$.
V Case II: Prime Power Dimensions
While prime dimensions provide a rigid structure where the Clifford group is maximal and irreducible, prime power dimensions $d=p^{m}$ (with $m\geq 2$) introduce a rich internal structure derived from the arithmetic of the ring $\mathbb{Z}/{p^{m}}\mathbb{Z}$. In this regime, the two primary pillars of universality no longer hold: the adjoint action becomes reducible, and the group loses its maximality.
V.1 Reducibility and Orbit Decomposition
In Sec. IV, we established that the irreducibility of the Clifford group’s action is equivalent to the transitivity of the $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ group on the Pauli indices (Lemma 15). While the prime case ($d=p$) guarantees transitivity, the prime power case ($d=p^{m}$, $m\geq 2$) introduces structural modifications due to the existence of non-trivial ideals in the ring $\mathbb{Z}/{p^{m}}\mathbb{Z}$.
Applying the machinery established in Lemma 15 and Lemma 16, we identify that the action of $\mathrm{SL}(2,\mathbb{Z}_{p^{m}})$ preserves the valuation of the index vector. Specifically, the greatest common divisor, $\gcd(u,v,p^{m})$, is an invariant of the $\mathrm{SL}(2,\mathbb{Z}_{p^{m}})$ action. Consequently, the set of Pauli indices partitions into $m$ distinct orbits.
By Lemma 15, each distinct orbit induces a distinct irreducible subrepresentation of the adjoint action. This leads to the following result.
Theorem 27.
For a prime power dimension $d=p^{m}$ ($m\geq 2$), the adjoint representation of the Clifford group on $\mathfrak{sl}(d,\mathbb{C})$ is reducible. The algebra decomposes into a direct sum of $m$ irreducible orthogonal invariant subspaces:
$\mathfrak{sl}(d,\mathbb{C})=\bigoplus_{k=0}^{m-1}\mathcal{W}_{k}.$ (18)
Each subspace $\mathcal{W}_{k}$ is spanned by the Pauli operators whose indices have a constant valuation:
$\mathcal{W}_{k}=\operatorname{span}\left\{P_{(u,v)}:\gcd(u,v,p^{m})=p^{k}\right\}.$ (19)
This result fundamentally shifts the criteria for universality compared to the prime case. Since the Lie algebra is fragmented into irreducible subspaces $\mathcal{W}_{k}$ that the Clifford group cannot connect, a supplementary gate must do more than simply generate infinite order; it must actively “mix” these subspaces, bridging irreducibility.
V.2 Orbit-mixing Gates and Irreducibility
To restore universality, we require a gate capable of coupling the orthogonal invariant subspaces $\mathcal{W}_{k}$ identified in Theorem 27. Since the Clifford group acts transitively only within each symplectic orbit, the extending gate must provide a mapping between them. We verify this capability by analyzing the Fourier spectrum of the gate’s adjoint action.
Consider the diagonal unitary gate $T_{\zeta}$ associated with a phase function $\zeta:\mathbb{Z}_{d}\to\mathrm{U}(1)$, as defined in Sec. IV.4. The adjoint action on an operator $X^{a}$ is given by:
$\operatorname{Ad}_{T_{\zeta}}(X^{a})=\zeta(a)X^{a}D_{a},$ (20)
where $D_{a}$ is the diagonal operator with entries $(D_{a})_{xx}=\delta_{\zeta}(a,x)$. Here, $\delta_{\zeta}$ is the differential previously introduced in Eq. (14).
To identify which subspaces $\mathcal{W}_{k}$ are populated by this operation, we expand the diagonal operator $D_{a}$ in the Pauli-$Z$ basis using the Discrete Fourier Transform (DFT). Recall that for a function $f:\mathbb{Z}_{d}\to\mathbb{C}$, the Fourier transform is defined as:
$\hat{f}(y)=\frac{1}{\sqrt{d}}\sum_{x\in\mathbb{Z}_{d}}f(x)\omega^{-xy}.$ (21)
Substituting the inverse DFT of the differential row $x\mapsto\delta_{\zeta}(a,x)$, the adjoint action expands in terms of the generalized Pauli basis $V_{(a,b)}=X^{a}Z^{b}$ as follows:
$\operatorname{Ad}_{T_{\zeta}}(X^{a})=\frac{\zeta(a)}{\sqrt{d}}\sum_{b\in\mathbb{Z}_{d}}\widehat{(\delta_{\zeta,a})}(b)V_{(a,b)}.$ (22)
This expansion reveals the selection rules of the gate. Recall that a basis vector $V_{(a,b)}$ belongs to the subspace $\mathcal{W}_{k}$ if the valuation of its indices is $\gcd(a,b,p^{m})=p^{k}$. Since diagonal gates preserve the row index $a$, a vector starting in a singular subspace $\mathcal{W}_{k}$ (where $\gcd(a,d)=p^{k}$ with $k\geq 1$) is constrained to that row. To map this vector into the generic subspace $\mathcal{W}_{0}$ (where the indices are units), the gate must generate a $Z$-component $b$ such that $\gcd(a,b)$ becomes a unit. This is algebraically possible if and only if $b$ itself is a unit.
Thus, irreducibility requires the gate to map at least one operator from every singular subspace into the generic subspace.
Definition 28.
A diagonal gate $T_{\zeta}$ is called an orbit-mixing gate if, for every proper divisor $u\in\{p,p^{2},\dots,p^{m-1}\}$, there exists at least one $b\in\mathbb{Z}_{d}^{\times}$ such that:
$\widehat{(\delta_{\zeta,u})}(b)\neq 0.$ (23)
Theorem 29.
Let $d=p^{m}$. If $T_{\zeta}$ is an orbit-mixing gate, then the adjoint representation of the extended group $G=\langle\mathcal{C}_{d},T_{\zeta}\rangle$ on $\mathfrak{sl}(d,\mathbb{C})$ is irreducible.
Proof.
Assume, for the sake of contradiction, that the representation is reducible. Let $\mathcal{S}\subset\mathfrak{sl}(d,\mathbb{C})$ be a proper, non-zero $G$-invariant subspace. Since the representation is unitary, $\mathcal{S}^{\perp}$ is also invariant. Because $\mathcal{C}_{d}\subset G$, any invariant subspace must be a direct sum of the Clifford-irreducible components $\mathcal{W}_{k}$.
We prove that any non-zero $G$-invariant subspace must contain the generic subspace $\mathcal{W}_{0}$. Let $\mathcal{S}$ be such a subspace.
  • Case 1: $\mathcal{S}$ contains a subspace $\mathcal{W}_{k}$ with $k\geq 1$. Let $X^{p^{k}}\in\mathcal{W}_{k}\subseteq\mathcal{S}$. Since $\mathcal{S}$ is $G$-invariant, $\operatorname{Ad}_{T_{\zeta}}(X^{p^{k}})\in\mathcal{S}$. By Eq. (22) and the Orbit-Mixing Condition (Def. 28), the expansion of this vector contains a term $\beta V_{(p^{k},b)}$ with $\beta\neq 0$ and $b\in\mathbb{Z}_{d}^{\times}$. Since $b$ is a unit, $\gcd(p^{k},b,p^{m})=1$, which implies $V_{(p^{k},b)}\in\mathcal{W}_{0}$. Since $\mathcal{S}$ is invariant under the Clifford group (which acts transitively on the orbit associated with $\mathcal{W}_{0}$), containing one vector in $\mathcal{W}_{0}$ implies $\mathcal{W}_{0}\subseteq\mathcal{S}$.
  • Case 2: $\mathcal{S}$ is contained entirely within $\mathcal{W}_{0}$. Since $\mathcal{S}$ is non-zero and decomposes into Clifford orbits, and $\mathcal{W}_{0}$ corresponds to a single transitive orbit, it must be that $\mathcal{S}=\mathcal{W}_{0}$.
In both cases, $\mathcal{W}_{0}\subseteq\mathcal{S}$. By the same logic, $\mathcal{W}_{0}\subseteq\mathcal{S}^{\perp}$. Thus, $\mathcal{W}_{0}\subseteq\mathcal{S}\cap\mathcal{S}^{\perp}=\{0\}$, a contradiction. ∎
V.3 Universality via the $T_{s}$-gate Family
We now apply our general machinery—irreducibility via orbit-mixing gates and infiniteness via the geometric criterion—to the generic family of diagonal gates $T_{s}$. Recall that $T_{s}$ introduces phases proportional to the roots of unity of order $s$, see Eq. (15).
Theorem 30.
Let $d=p^{m}$ with $m\geq 2$. Consider the group $G_{s}=\langle\mathcal{C}_{d},T_{s}\rangle$ generated by the Clifford group and the gate $T_{s}$ (Eq. (15)).
  1. 1.
    The group $G_{s}$ acts irreducibly on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$ if and only if $s\nmid d$.
  2. 2.
    The group $G_{s}$ is single-qudit universal (dense in $\mathrm{SU}(d)$) if $s\nmid d$ and $s$ satisfies the bound:
    $s>\frac{\pi(d-1)}{2\arcsin(1/4)}.$ (24)
Proof.
1. Proof of Irreducibility. We verify that $T_{s}$ satisfies Definition 28 to apply Theorem 29. The phase function is $\zeta(k)=e^{2\pi ik/s}$, and the associated 1-coboundary $\delta_{\zeta}(u,k)=\zeta(u+k)/\zeta(u)\zeta(k)$ has the form:
$\delta_{\zeta}(u,k)=\begin{cases}1&\text{if }u+k<d,\\ \lambda_{s}&\text{if }u+k\geq d,\end{cases}$ (25)
where $\lambda_{s}=\exp(-2\pi id/s)$.
We must show that for any proper divisor $u$ of $d$, the DFT of the function $f_{u}(k)=\delta_{\zeta}(u,k)$ has full support on the units of $\mathbb{Z}_{d}$. The DFT at frequency $n$ is given by:
$\sqrt{d}\hat{f}_{u}(n)=\sum_{k=0}^{d-1}f_{u}(k)\omega^{-nk}=\sum_{k=0}^{d-u-1}\omega^{-nk}+\lambda_{s}\sum_{k=d-u}^{d-1}\omega^{-nk},$
where $\omega=e^{2\pi i/d}$. For any unit $n\in\mathbb{Z}_{d}^{\times}$, we have $\omega^{-n}\neq 1$. Summing the geometric series for the two intervals yields:
$\sqrt{d}\hat{f}_{u}(n)$ $=\frac{1-\omega^{-n(d-u)}}{1-\omega^{-n}}+\lambda_{s}\frac{\omega^{-n(d-u)}-\omega^{-nd}}{1-\omega^{-n}}$
$=\frac{1-\omega^{nu}+\lambda_{s}(\omega^{nu}-1)}{1-\omega^{-n}}\quad(\text{since }\omega^{-nd}=1)$
$=\frac{(1-\omega^{nu})(1-\lambda_{s})}{1-\omega^{-n}}.$
We examine the factors in the numerator to determine if the spectrum vanishes:
  • The term $(1-\omega^{nu})$ is non-zero because $n$ is a unit (coprime to $d$) and $u$ is a proper divisor ($0<u<d$), so $nu$ is not a multiple of $d$.
  • The term $(1-\lambda_{s})$ vanishes if and only if $\lambda_{s}=1$. Recalling that $\lambda_{s}=e^{-2\pi id/s}$, this occurs strictly when $d/s$ is an integer.
Thus, $\hat{f}_{u}(n)\neq 0$ for all units $n$ if and only if $s\nmid d$. This confirms that the condition $s\nmid d$ is necessary and sufficient for irreducibility.
2. Proof of Universality. According to Theorem 4, universality requires both irreducibility (established in Item 1) and infiniteness. To prove infiniteness, we apply Theorem 10. The eigenvalues of $T_{s}$ are $\{1,e^{2\pi i/s},\dots,e^{2\pi i(d-1)/s}\}$. The spectral span is exactly:
$\ell(T_{s})=\frac{d-1}{s}.$ (26)
From Corollary 14, the group is guaranteed to be infinite if the projective distance associated with this span is less than $1/2$. The precise condition is:
$\pi\ell(T_{s})<2\arcsin(1/4).$ (27)
Substituting the expression for $\ell(T_{s})$ and solving for $s$:
$\frac{\pi(d-1)}{s}<2\arcsin(1/4)\implies s>\frac{\pi(d-1)}{2\arcsin(1/4)}.$
If this bound holds, the group is infinite. Combined with the irreducibility from Item 1, the group is dense in $\mathrm{SU}(d)$. ∎
VI Case III: Composite Dimensions
We now turn to the final regime: composite dimensions $d=d_{1}d_{2}\cdots d_{n}$ where the factors are pairwise coprime. Unlike prime or prime-power dimensions, which typically require the direct injection of diagonal non-Clifford gates (such as the qubit $T$-gate or the qudit $T_{s}$), composite dimensions offer a path to universality driven by “structural mixing.”
In this regime, the necessary resources are not diagonal non-Clifford gates, but rather standard entangling gates (generalized CNOTs) that couple different coprime factors: intra-qudit gates. While these coupling gates are external to the local Clifford product, they are “natural” resources in any register-based architecture: they are permutation matrices that assist entanglement generation. The central result of this Section is the demonstration that the arithmetic interplay between coprime dimensions, coupled by intra-qudit gates, spontaneously generates the non-Clifford diagonal resources required for universality.
VI.1 Irreducibility via Tensor Mixing
To establish universality in composite dimensions $d=kl$, we first determine the conditions under which a group generated by local and entangling gates acts irreducibly on the global Lie algebra. Since a necessary condition for the density of a subgroup in $\mathrm{SU}(d)$ is the irreducibility of its action on the complexified Lie algebra $\mathfrak{sl}(d,\mathbb{C})$, as discussed in Sec. II, we frame our analysis in $\mathfrak{sl}(d,\mathbb{C})$.
We begin by decomposing the global algebra under the action of the local subgroup $\mathrm{SU}(k)\otimes\mathrm{SU}(l)$.
Proposition 31.
Under the adjoint action of the product group $\mathrm{SU}(k)\otimes\mathrm{SU}(l)$, the Lie algebra $\mathfrak{sl}(kl,\mathbb{C})$ decomposes into three orthogonal irreducible representations:
$\begin{split}\mathfrak{sl}(kl,\mathbb{C})\cong&\underbrace{(\mathfrak{sl}(k,\mathbb{C})\otimes I_{l})}_{\mathfrak{p}_{1}}\oplus\underbrace{(I_{k}\otimes\mathfrak{sl}(l,\mathbb{C}))}_{\mathfrak{p}_{2}}\\ &\oplus\underbrace{(\mathfrak{sl}(k,\mathbb{C})\otimes\mathfrak{sl}(l,\mathbb{C}))}_{\mathfrak{m}}.\end{split}$ (28)
Here, $\mathfrak{h}=\mathfrak{p}_{1}\oplus\mathfrak{p}_{2}$ represents the subalgebra of local operations.
Proof.
Consider the isomorphism $M_{kl}(\mathbb{C})\cong M_{k}(\mathbb{C})\otimes M_{l}(\mathbb{C})$. By applying the decomposition of the general linear algebra into scalar and traceless parts, $M_{n}(\mathbb{C})=\mathbb{C}I_{n}\oplus\mathfrak{sl}(n,\mathbb{C})$, we expand the tensor product as follows:
$M_{kl}(\mathbb{C})$ $\cong\big(\mathbb{C}I_{k}\oplus\mathfrak{sl}(k,\mathbb{C})\big)\otimes\big(\mathbb{C}I_{l}\oplus\mathfrak{sl}(l,\mathbb{C})\big)$
$\cong(\mathbb{C}I_{k}\otimes\mathbb{C}I_{l})\oplus(\mathfrak{sl}(k,\mathbb{C})\otimes I_{l})$
$\quad\oplus(I_{k}\otimes\mathfrak{sl}(l,\mathbb{C}))\oplus(\mathfrak{sl}(k,\mathbb{C})\otimes\mathfrak{sl}(l,\mathbb{C})).$
The first term simplifies to $\mathbb{C}(I_{k}\otimes I_{l})=\mathbb{C}I_{kl}$, which corresponds to the global scalar matrices. Consequently, the direct sum of the remaining three terms constitutes the trace-zero subalgebra $\mathfrak{sl}(kl,\mathbb{C})$.
[…]
To analyze the irreducibility of the extended group $G=\langle\mathrm{SU}(k)\otimes\mathrm{SU}(l),V\rangle$, where $V\in\mathrm{SU}(kl)$ is an intra-qudit gate, we must determine whether the decomposition in Eq. (28) remains invariant under the full group. Observe that the subalgebra of local operations, $\mathfrak{h}=\mathfrak{p}_{1}\oplus\mathfrak{p}_{2}$, is invariant under the local group by definition. If the gate $V$ also maps $\mathfrak{h}$ to itself (i.e., if $V$ normalizes the local algebra), then $\mathfrak{h}$ becomes a proper $G$-invariant subspace, preventing the global action from being irreducible. Consequently, identifying the subgroup of $\mathrm{SU}(kl)$ that preserves this local structure is a necessary prerequisite for establishing universality.
Lemma 32.
Let $H=\mathrm{SU}(k)\otimes\mathrm{SU}(l)$ with integers $k,l\geq 2$. The normalizer of $H$ in the global group $\mathrm{SU}(kl)$ is:
  • $H$ itself, if $k\neq l$.
  • The group generated by $H$ and the SWAP operator $S$, if $k=l$, where $S(u\otimes v)=v\otimes u$.
Proof.
Let $\mathfrak{h}=\mathfrak{p}_{1}\oplus\mathfrak{p}_{2}$ be the Lie algebra of $H$. We determine the normalizer in the algebra $\mathfrak{g}=\mathfrak{sl}(kl,\mathbb{C})$. Let $X\in\mathfrak{g}$ be an element such that $X,\mathfrak{h}]\subseteq\mathfrak{h}$. Decomposing $X=X_{\mathfrak{h}}+X_{\mathfrak{m}}$ according to Eq. (28), the condition implies $X_{\mathfrak{m}},\mathfrak{h}]\subseteq\mathfrak{h}$. However, $\mathfrak{m}$ is an invariant subspace under $\mathfrak{h}$ (it is a representation of $H$), so $X_{\mathfrak{m}},\mathfrak{h}]\subseteq\mathfrak{m}$. Since $\mathfrak{h}\cap\mathfrak{m}=\{0\}$, the bracket must vanish: $X_{\mathfrak{m}},\mathfrak{h}]=0$. Since, $\mathfrak{m}\cong\mathfrak{sl}(k,\mathbb{C})\otimes\mathfrak{sl}(l,\mathbb{C})$ is a non-trivial irreducible representation of $\mathfrak{h}$; by Schur’s Lemma, the only element that commutes with the entire action is zero. Thus, $X_{\mathfrak{m}}=0$, implying the Lie algebra normalizer is exactly $\mathfrak{h}$.
At the group level, any normalizer element $V\notin H$ must induce an outer automorphism of the ideal structure of $\mathfrak{h}=\mathfrak{su}(k)\oplus\mathfrak{su}(l)$.
  • If $k\neq l$, the ideals have different dimensions and cannot be permuted. Since $\mathrm{SU}(n)$ has no outer automorphisms realized by unitary conjugation (which preserves spectrum), $V$ must be in $H$.
  • If $k=l$, the ideals are isomorphic, and the SWAP operator $S$ permutes them. Thus, the normalizer is the extension of $H$ by $S$.
We now state the main sufficient condition for irreducibility in composite systems.
Theorem 33.
Let $k,l\geq 2$. Let $G_{k}\subseteq\mathrm{SU}(k)$ and $G_{l}\subseteq\mathrm{SU}(l)$ be subgroups acting irreducibly on $\mathfrak{sl}(k,\mathbb{C})$ and $\mathfrak{sl}(l,\mathbb{C})$, respectively. Let $V\in\mathrm{SU}(kl)$ be a unitary gate. The group $G=\langle G_{k}\otimes I,I\otimes G_{l},V\rangle$ acts irreducibly on $\mathfrak{sl}(kl,\mathbb{C})$ via the adjoint representation if and only if:
  1. 1.
    $V\notin\mathrm{SU}(k)\otimes\mathrm{SU}(l)$, for $k\neq l$.
  2. 2.
    $V\notin\langle\mathrm{SU}(k)^{\otimes 2},S\rangle$, for $k=l$.
Proof.
If $V$ belongs to the normalizer described in Lemma 32, then the product subalgebra $\mathfrak{h}=\mathfrak{p}_{1}\oplus\mathfrak{p}_{2}$ remains invariant under the full group $G$. Since $\mathfrak{h}\subsetneq\mathfrak{sl}(kl,\mathbb{C})$, the action is reducible.
Now, suppose $V$ satisfies the conditions (i.e., $V$ does not normalize the product structure). Let $W\subseteq\mathfrak{sl}(kl,\mathbb{C})$ be a non-zero subspace invariant under $G$. We prove $W=\mathfrak{sl}(kl,\mathbb{C})$.
Since the local subgroup $H_{0}=G_{k}\otimes G_{l}$ is contained in $G$, $W$ must be a direct sum of the irreducible $H_{0}$-modules defined in Proposition 31: $\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{m}$. The unitary adjoint action preserves the inner product, so the orthogonal complement $W^{\perp}$ is also a $G$-invariant subspace and decomposes similarly. This symmetry implies that if $W\neq\mathfrak{sl}(kl,\mathbb{C})$, then either $W$ or $W^{\perp}$ must be contained entirely within the product subalgebra $\mathfrak{h}=\mathfrak{p}_{1}\oplus\mathfrak{p}_{2}$. Let $S\subseteq\mathfrak{h}$ be such a non-zero proper invariant subspace. We analyze two cases:
  • If $S=\mathfrak{h}$, then $\operatorname{Ad}_{V}(\mathfrak{h})=\mathfrak{h}$. This implies $V$ normalizes the product subalgebra, which is excluded by hypothesis.
  • If $S=\mathfrak{p}_{1}$ (without loss of generality), then $V$ maps $\mathfrak{p}_{1}$ to itself. Note that $\mathfrak{p}_{2}$ corresponds exactly to the centralizer of $\mathfrak{p}_{1}$ in the algebra, i.e., $\{Y\in\mathfrak{sl}(kl):[Y,\mathfrak{p}_{1}]=0\}=\mathfrak{p}_{2}$. Since conjugation by $V$ is a Lie algebra automorphism, it maps the centralizer of a set to the centralizer of its image. Thus, $V$ must also preserve $\mathfrak{p}_{2}$. Consequently, $V$ preserves $\mathfrak{p}_{1}\oplus\mathfrak{p}_{2}=\mathfrak{h}$, leading back to the first contradiction.
We conclude that no proper subspace of the algebra can be invariant. Thus, $W$ must contain the correlation space $\mathfrak{m}$ and the local spaces, implying $W=\mathfrak{sl}(kl,\mathbb{C})$. ∎
Combining Theorem 33 and Theorem 4, we derive sufficient conditions for universality in composite systems.
Corollary 34.
Let $G_{k}\subset\mathrm{SU}(k)$ and $G_{l}\subset\mathrm{SU}(l)$ be subgroups acting irreducibly on $\mathfrak{sl}(k,\mathbb{C})$ and $\mathfrak{sl}(l,\mathbb{C})$, respectively. Let $V\in\mathrm{SU}(kl)$. The group $G=\langle G_{k}\otimes I,I\otimes G_{l},V\rangle$ is universal if and only if the following two conditions hold:
  1. 1.
    $V\notin N_{\mathrm{SU}(kl)}(\mathrm{SU}(k)\otimes\mathrm{SU}(l))$,
  2. 2.
    The group $G$ is infinite.
As a particular case we obtain:
Corollary 35.
Let $k,l\geq 2$. Suppose $\mathcal{S}_{k}\subset\mathrm{SU}(k)$ and $\mathcal{S}_{l}\subset\mathrm{SU}(l)$ are sets of gates that are single-qudit universal (i.e., $\overline{\langle\mathcal{S}_{k}\rangle}=\mathrm{SU}(k)$ and $\overline{\langle\mathcal{S}_{l}\rangle}=\mathrm{SU}(l)$). Let $V\in\mathrm{SU}(kl)$ be an additional gate. The set $\mathcal{S}=(\mathcal{S}_{k}\otimes\{I\})\cup(\{I\}\otimes\mathcal{S}_{l})\cup\{V\}$ is universal for quantum computation on the composite system (i.e., $\overline{\langle\mathcal{S}\rangle}=\mathrm{SU}(kl)$) if and only if:
  1. 1.
    If $k\neq l$, $V\notin\mathrm{SU}(k)\otimes\mathrm{SU}(l)$.
  2. 2.
    If $k=l$, $V\notin\mathrm{SU}(k)\otimes\mathrm{SU}(k)$ and $V\notin(\mathrm{SU}(k)\otimes\mathrm{SU}(k))\cdot S$.
This result generalizes the seminal theorem of Brylinski and Brylinski [Brylinski2002], which established that adjoining any entangling gate to the local unitary group $\mathrm{SU}(d)\otimes\mathrm{SU}(d)$ yields a universal set for two-qudit quantum computation.
VI.2 Intra-qudit Gates and Induced Magic
To implement the tensor mixing strategy, we employ the generalized intra-qudit CNOT gate. This gate serves as the fundamental entangling resource for qudit registers of arbitrary dimensions.
Definition 36.
[…]
This gate is a permutation matrix that is always unitary, regardless of the relationship between $p$ and $q$. However, its interaction with local Pauli operators produces a phase kickback effect that depends strictly on the arithmetic properties of the dimensions.
Lemma 37.
Conjugating the inverse local Pauli operator $I_{p}\otimes Z_{q}^{\dagger}$ by the intra-qudit gate yields a separable operator containing the generic phase gate $T_{q}$:
$CN_{p,q}(I_{p}\otimes Z_{q}^{\dagger})CN_{p,q}^{\dagger}=T_{q}\otimes Z_{q}^{\dagger}.$ (30)
Proof.
[…]
The resulting operator is the tensor product of $|T_{q}=\sum\omega^{x}\ket{x}\bra{x}$ acting on the control qudit and the local Pauli $Z_{q}^{\dagger}$ acting on the target. ∎
The induced resource $T_{q}$ derived in Lemma 37 injects a diagonal gate into the Hilbert space of dimension $p$. By using $T_{q}$ with the local Clifford resources (specifically the Pauli $Z$, which corresponds to $T_{p}$), we can construct a local $T_{pq}$ in the space of dimension $p$.
Proposition 38.
Let $p$ and $q$ be coprime integers. Let $a,b\in\mathbb{Z}$ be the coefficients satisfying the Bézout identity $ap+bq=1$. By combining the local Pauli operator $T_{p}$ and the induced resource $T_{q}$, we construct the generic phase gate of order $pq$:
$|.T_{pq}=T_{q}^{a}T_{p}^{b}=\sum_{k=0}^{p-1}\exp\left(\frac{2\pi ik}{pq}\right)\ket{k}\bra{k}$. (31)
Proof.
[…]
$\lambda_{k}=\left(e^{2\pi ik/q}\right)^{a}\cdot\left(e^{2\pi ik/p}\right)^{b}=\exp\left(2\pi ik\left[\frac{a}{q}+\frac{b}{p}\right]\right).$
Finding a common denominator and applying the Bézout identity $ap+bq=1$, the exponent simplifies to
$\frac{a}{q}+\frac{b}{p}=\frac{ap+bq}{pq}=\frac{1}{pq}.$
Substituting this back, we obtain $\lambda_{k}=e^{2\pi ik/pq}$, which exactly matches the definition of the gate $T_{pq}$. ∎
VI.3 Main Theorem: Universality
We now assemble the components—global irreducibility via Theorem 33 and local infiniteness via the $T$’s constructed—to prove the central result of this work.
Theorem 39.
Let $d=d_{1}d_{2}\cdots d_{n}$ be a proper composite dimension decomposed into pairwise coprime prime power factors (i.e., $d_{k}=p_{k}^{m_{k}}$ with $\gcd(d_{i},d_{j})=1$ for $i\neq j$). The group generated by the single-qudit Clifford group and the intra-qudit gates; that is, the group
$G=\langle\mathcal{C}_{d},\{CN_{k,k+1}\}_{k=1}^{n-1}\rangle,$
is dense in $\mathrm{SU}(d)$.
Proof.
We verify the conditions of Corollary 34.
1. Establishment of Local Irreducibility. First, we ensure that the local action on every factor is irreducible. Consider a subsystem $k$ of dimension $d_{k}$.
  • If $d_{k}=p$ is prime, by Theorem 17, the local Clifford group $\mathcal{C}_{d_{k}}$ acts irreducibly on $\mathfrak{sl}(d_{k},\mathbb{C})$.
  • If $d_{k}=p^{m}$ is a prime power ($m\geq 2$), using Lemma 37, we obtain a local gate $T_{d_{j}}$. Since $\gcd(d_{k},d_{j})=1$, the order $s=d_{j}$ satisfies $s\nmid d_{k}$. Thus, $T_{d_{j}}$ satisfies the condition of Item 1 in Theorem 30, and consequently the local group $\langle\mathcal{C}_{d_{k}},T_{d_{j}}\rangle$ acts irreducibly.
With irreducible local algebras established, the gates $\{CN_{k,k+1}\}$ serve as the mixing operators $V$. They do not normalize the product structure. By applying Theorem 33, the global group $G$ acts irreducibly on the global algebra $\mathfrak{sl}(d,\mathbb{C})$.
2. Verification of Infiniteness. To prove universality, it remains only to show that $G$ is infinite. Since global irreducibility is guaranteed, it suffices to demonstrate that the group contains a dense subgroup on at least one local subsystem. We identify such a subsystem by distinguishing two exhaustive cases based on the nature of the factorization of $d$:
  • Case A: The factorization contains at least one prime factor. Suppose there exists a subsystem $k$ such that $d_{k}=p$ is prime. Let $d_{j}$ be its coprime neighbor. We verify density on subsystem $k$. Using Proposition 38, a local gate $T_{S}$ with $S=d_{k}d_{j}$ is constructed. Since $\gcd(d_{k},d_{j})=1$, we have $S\nmid d_{k}$ (and thus $S\nmid K_{d_{k}}$). By Corollary 26, the extended local group is dense in $\mathrm{SU}(p)$. Thus, the global group is infinite.
  • Case B: The factorization consists entirely of composite prime powers. Suppose all factors are proper prime powers ($d_{k}=p_{k}^{m_{k}}$ with $m_{k}\geq 2$). By Theorem 30, the group is dense on factor $k$ if $S=d_{k}d_{j}$ satisfies the bound:
    $S>\frac{\pi(d_{k}-1)}{2\arcsin(1/4)}.$
    Consider the worst-case scenario with the smallest possible coprime proper powers: $d_{k}=4$ and $d_{j}=9$. The bound for $d_{k}=4$ is $S>6.22(3)\approx 18.66$ and $S=4\times 9=36$. Since $36>18.66$, the condition is strictly satisfied. Any other pair of proper powers yields a larger product, satisfying the bound even more robustly. Therefore, the group is dense in the subsystem $k$.
In both scenarios, we guarantee the existence of at least one locally dense subsystem. Combined with the global irreducibility established in the first part, this implies that $G$ is dense in the full group $\mathrm{SU}(d)$. ∎
VI.4 Explicit Universal Gate Sets
A consequence of Theorem 39 is that for composite systems with coprime dimensions, universality is achieved using only standard architectural primitives. In the qubit regime ($d=2$), the group generated by local Clifford gates and CNOTs remains finite (the global Clifford group), requiring the injection of a dedicated magic resource (such as a $T$-gate) to achieve density. In contrast, for coprime composite systems, the inclusion of the intra-qudit gate CNCN—typically considered a standard permutation—is sufficient to break the finite group structure. Due to the arithmetic mismatch between factors, the interplay between the local Clifford groups and the intra-qudit gate generates an infinite, dense group without requiring explicit diagonal magic gates.
Corollary 40.
Let $d=d_{1}d_{2}\cdots d_{n}$ be a proper composite dimension with pairwise coprime prime power factors. The following finite set of gates constitutes a universal set for quantum computation on arrays of registers of dimension $d$:
$\mathcal{G}=\{X,H_{d},P\}\cup\bigcup_{k=1}^{n-1}\{CN_{d_{k},d_{k+1}}\}\cup\{CN_{d,d}\}.$ (32)
Remark 41.
The set $\mathcal{G}$ satisfies the universality condition of Theorem 2. Note the distinction between the entangling resources: $CN_{d_{k},d_{k+1}}\in\mathrm{SU}(d)$ is an intra-qudit operation that couples internal coprime factors and acts as the generator of non-Clifford resources. In contrast, $CN_{d,d}\in\mathrm{SU}(d^{2})$ represents the canonical inter-qudit gate acting on two logical qudits.
VII Discussion and Outlook
The results of this work show that the resources required to supplement the single-qudit Clifford group for universal quantum computation depend fundamentally on the arithmetic structure of the Hilbert space dimension. This dependence is not merely quantitative—involving more or fewer extra gates—but also qualitative: the type of resource changes across the trichotomy. This establishes dimensionality as a computational resource: the factorization properties of the dimension influence not only the cost but the nature of universality.
Our perspective reframes the universality question. Rather than asking which gates generate a dense subgroup of $\mathrm{SU}(d)$, we ask which operations compensate for structural limitations of the Clifford group. Gates that are merely non-Clifford never suffice unless $d=p$, where non-Cliffordness implies magic. Only those that have genuinely “magic” guarantee single-qudit universality. In composite dimensions, these notions diverge: adding a non-Clifford gate can enlarge the group without making it infinite, leaving universality unattained. Achieving universality, therefore, requires gates that both escape the Clifford group and ensure infiniteness. The failure of density may arise from reducibility of the adjoint action, finiteness of the group, or both—and the prime factorization of the dimension dictates which condition applies.
The classification we have presented reveals an unexpected connection between number theory and quantum computation. The prime factorization of the Hilbert space dimension determines whether universality requires explicit magic injection or emerges from structure. This insight highlights dimensionality as a design lever: by choosing coprime or prime-power architectures, one can trade off magic resources against control complexity. This connection is not merely formal; it may have concrete implications for the design of quantum processors.
The broader lesson is that quantum resources are context-dependent. Entanglement, magic, and coherence are necessary for quantum advantage, but which resources are “free” and which must be engineered depends on the computational framework. Dimensionality is context-dependent as well. Its arithmetic structure governs whether magic is intrinsic or must be injected. In homogeneous qubit architectures, entangling gates are free but magic is costly. In coprime architectures, both are free—the price to pay lies in the complexity of controlling heterogeneous subsystems. Understanding these trade-offs across different physical platforms is essential for identifying the most efficient routes to practical quantum computation.
Several limitations and open questions remain. For instance, multi-qudit universality, i.e. density in $\mathrm{SU}(d^{n})$, additionally requires an entangling gate satisfying the Brylinski criterion [Brylinski2002]. For coprime architectures, intra-qudit CNOT gates can serve as both magic and entangling resources, though compilation overhead in heterogeneous registers remains an open question. Also, realizing coprime dimensions physically poses challenges: high-fidelity gates between subsystems of different dimensions, such as a generalized intra-qudit CNOT between a qutrit and a ququart, demand precise control of a twelve-dimensional Hilbert space.
On the other hand, resource-theoretic measures of magic-such as stabilizer nullity [Beverland_2020], mana [Veitch2014], robustness [Howard2017], and stabilizer extent [Bravyi_2019] must be understood when magic emerges from coupling rather than injection. Beyond qudits, whether analogues of our trichotomy exist in continuous-variable systems, which have their own Clifford group and magic resources [Bartlett2002], or in hybrid architectures, is an intriguing question. Additionally, the complexity-theoretic implications of universality in the composite case remain unclear: does adding coprime Clifford + inter-qudit gates interpolate between classically simulable and universal circuits in a meaningful way?
Current architectures typically use homogeneous registers—arrays of identical qubits or qudits—where the Clifford group plus inter-qudit entangling gates remains finite and classically simulable via stabilizer methods [Gottesman1998, AaronsonGottesman2004, Campbell2012], so universality requires magic state distillation [Bravyi2005]. We propose an alternative: heterogeneous coprime architectures, where subsystems of dimensions $d_{1}$ and $d_{2}$ with $\gcd(d_{1},d_{2})=1$ achieve universality using only local Clifford gates and native inter-qudit, coupling. This principle could be realized in trapped ions [RingbauerEtAl2022, Low2020], superconducting circuits [Chicco_2023, Morvan2021], and photonic systems [Wang2020review, liu2024]. The main advantage is eliminating magic state distillation overhead, which dominates resource costs in standard fault-tolerant schemes [Litinski2019, Gidney_2021], even with recent improvements [PrakashSaha2025]. A potential drawback is increased control complexity for heterogeneous registers, but similar challenges already arise in modular and hybrid architectures.
Finally, we note that the trichotomy discovered here may be the first instance of a broader phenomenon. Whenever a mathematical structure admits a notion of “dimension” with nontrivial factorization properties, similar classifications may arise. Exploring such analogies-in continuous-variable systems, in topological quantum computation, or in other algebraic settings-offers an interesting direction for future work.
Acknowledgements.
C.G. was partially supported by Grant INV-2025-213-3452 from the School of Science of Universidad de los Andes. J.R. acknowledges support from the Office of the Vice-president of Research and Creative Activities and the Office of the Faculty of Science Vice-president of Research of Universidad de los Andes under the FAPA grant.

### v2

@@
 Abstract
-The efficient classical simulation of Clifford circuits constitutes a fundamental barrier to quantum advantage, typically overcome by injecting explicit non-Clifford “magic” resources.
-We demonstrate that for high-dimensional quantum systems (qudits), the resources required to break this barrier are strictly governed by the number-theoretic structure of the Hilbert space dimension $d$.
-By analyzing the adjoint action of the Clifford group, we establish a classification of single-qudit universality as a trichotomy.
-(I) For prime dimensions, the Clifford group is a maximal finite subgroup, and universality is robustly achieved by any non-Clifford gate.
-(II) For prime-power dimensions, the group structure fragments, requiring tailored diagonal non-Clifford gates to restore irreducibility.
-(III) Most notably, for composite dimensions with coprime factors, we demonstrate that standard entangling operations alone—specifically, generalized intra-qudit CNOT gates—generate the necessary non-Clifford resources to guarantee a dense subgroup of $\mathrm{SU}(d)$ without explicit diagonal magic injection.
-Our proofs rely on a new geometric criterion establishing that a subgroup with irreducible adjoint action is infinite if it contains a non-scalar element with projective distance strictly less than $1/2$ from the identity.
-These results suggest that “coprime architectures”—hybrid registers combining incommensurate dimensions—can bypass the standard overhead of magic state distillation.
+The Clifford group is efficiently classically simulable, and universality is obtained by supplementing it with non-Clifford resources.
+We determine which single-qudit gates suffice to achieve universality.
+We show that the structure of such resources is governed by the prime factorization of the qudit dimension $d$.
+Using the adjoint action on the space of complex trace-zero matrices, we relate density to irreducibility together with an infiniteness criterion, yielding a trichotomy based on the factorization of $d$.
+When $d$ is prime, any non-Clifford gate generates a dense subgroup of the determinant-one unitaries.
+If $d$ is a prime power, the adjoint action is reducible, and universality requires gates that couple the resulting invariant subspaces.
+For composite $d$ with pairwise coprime factors, generalized intra-qudit controlled-NOT gates connecting the factors already suffice.
+These findings suggest that “composite architectures”—hybrid registers combining incommensurate dimensions—offer a route to bypass the standard overhead associated with magic-state injection.
 I Introduction
-What makes a quantum computer more powerful than a classical one?
-The standard narrative points to interference and entanglement, yet the Gottesman-Knill theorem reveals that highly-entangled states can be efficiently simulated classically when generated by Clifford operations alone [Gottesman1998, AaronsonGottesman2004].
-The missing resource that lifts computation beyond classical reach is typically identified as “magic”: non-stabilizer states or non-Clifford gates that enable quantum universality—the ability to approximate any unitary operator to arbitrary precision by composing elements from a finite gate set [Kitaev1997, Deutsch1995].
-This paper reveals that for quantum systems of composite dimension magic can emerge from the interplay of classical operations, with the local Hilbert space dimension unfolding as a previously unrecognized computational resource.
-Consider a quantum register composed of subregisters with coprime local dimensions—say, a qutrit ($d=3$) coupled to a ququart ($d=4$).
-Coupling between them can be attained via the generalized intra-qudit controlled-addition gate: a permutation matrix that adds the control value to the target modulo $d$.
-Notably, this operation is classical and specified entirely by discrete modular arithmetic.
-Yet we prove that when such intra-qudit gates connect subregisters of coprime dimension, the group they generate together with local Clifford operations is dense in the special unitary group.
-Universality emerges from arithmetic alone.
-This inverts the conventional understanding of quantum resources.
-In standard qubit architectures, the single-qubit Clifford group supplemented by CNOT gates remains finite—it is the multi-qubit Clifford group, efficiently simulable by the Gottesman-Knill theorem.
-Achieving universality demands a dedicated non-Clifford resource, such as the $T$-gate or magic states prepared through costly distillation protocols [Bravyi2005].
-Our results show that this overhead is an artifact of homogeneous architectures built from identical subsystems, not an intrinsic feature of quantum computation.
-When local dimensions are coprime, the “free” entangling operations that any architecture provides are already sufficient.
-The phenomenon has a transparent algebraic origin.
-The generalized CNOT is a classical reversible gate.
-However, when connecting subsystems of coprime dimensions $p$ and $q$, it acts on local Pauli operators to produce diagonal gates.
-By Bézout’s identity, these gates construct fractional phases $2\pi/pq$—inaccessible to either local Clifford group.
-Hence, the generation of dense subgroups follows from gates that individually preserve the computational basis.
-This suggests viewing local dimension as a computational resource.
-Unlike prime-power dimensions like $d=4$, coprime dimensions like $d=6$ allow the “mixing” of incompatible modular arithmetics, creating quantum resources from the arithmetic structure alone.
-A finitely generated subgroup of $\mathrm{SU}(d)$ is dense if and only if it is infinite and acts irreducibly on the Lie algebra $\mathfrak{su}(d)$ via the adjoint representation.
-When applied to extensions of the Clifford group, this criterion identifies the specific obstructions to universality.
-In composite dimensions, unlike the prime case, the Clifford group is not maximal; therefore, a non-Clifford gate may generate a strictly larger yet still finite group.
-To rule out this possibility and ensure infiniteness, we introduce a geometric criterion: a subgroup with irreducible adjoint action is infinite if it contains a non-scalar element within projective distance $1/2$ of the identity.
+A central question in quantum information science is understanding the computational advantage of quantum computers over classical architectures.
+While quantum entanglement is often cited as the source of such an advantage, the Gottesman-Knill theorem reveals a fundamental limitation: highly-entangled states generated by Clifford operations alone can be efficiently simulated classically [19, 1].
+This result extends to higher-dimensional qudits, where efficient classical simulation of Clifford circuits has been established [15, 12].
+The missing resource that lifts computation beyond classical reach is typically identified as magic: non-stabilizer states or non-Clifford gates that enable quantum universality—the ability to approximate any unitary operator to arbitrary precision by finite gate compositions [24].
+Establishing conditions under which the Clifford group in high-dimensional systems yields universal quantum computation remains an open problem.
+In this work, we analyze this question for the single-qudit Clifford group.
+[…]
+Establishing universality amounts to proving that the generated group is dense in $\mathrm{SU}(d)$, the group of unitary operators with determinant one.
+We address this topological problem via the adjoint representation of the group on $\mathfrak{sl}(d,\mathbb{C})$.
+A finitely generated subgroup of $\mathrm{SU}(d)$ is dense if and only if it is infinite and acts irreducibly on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$ (Theorem 4).
+The extent to which the Clifford group satisfies these conditions is governed by the prime factorization $d=p_{1}^{m_{1}}\cdots p_{k}^{m_{k}}$ (Theorem 24 and Theorem 30).
+This factorization dictates the algebraic obstruction, establishing a natural trichotomy.
+When $d=p$ is prime, the Clifford group exhibits maximal rigidity.
+We prove that its adjoint action is irreducible if and only if $d$ is prime (Theorem 24).
+Furthermore, the group is projectively maximal if and only if $d$ is prime: any strictly larger finite subgroup must differ only by scalar phases (Theorem 25).
+Consequently, adding any non-Clifford gate forces universality (Corollary 27).
+However, when the factorization includes higher powers ($m_{i}\geq 2$) or multiple distinct primes ($k>1$), this rigidity breaks down.
+The adjoint action becomes reducible, fragmenting the Lie algebra into invariant subspaces indexed by divisibility properties of $d$ (Theorem 30).
+Achieving universality, therefore, requires resources capable of coupling these subspaces to restore irreducibility.
+[…]
+We establish density by verifying two conditions.
+First, irreducibility requires characterizing the invariant subspaces of $\mathfrak{sl}(d,\mathbb{C})$ under the Clifford group’s adjoint action.
+We classify these subspaces via the symplectic orbit structure of $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ and demonstrate that our resource gates mix the resulting components.
+Second, infiniteness is established through a simple criterion: a subgroup with irreducible adjoint action is infinite if it contains a non-scalar element within projective distance $1/2$ of the identity.
+This criterion reduces the verification to a computation over the eigenvalues.
 Related work.
-The structure of universality for qubit systems is well understood: the Clifford group supplemented by any non-Clifford gate—typically the $T$-gate from the third level of the Clifford hierarchy—generates a dense subgroup of $\mathrm{SU}(2)$ [Gottesman1999, Nebe2001].
-This “Clifford + TT” paradigm underlies modern fault-tolerant quantum computation, where magic states enable non-Clifford operations through gate teleportation [Bravyi2005, Knill2004].
+The structure of universality for qubit systems is well understood: the Clifford group supplemented by any non-Clifford gate—typically the $T$-gate from the third level of the Clifford hierarchy [18]—generates a dense subgroup of $\mathrm{SU}(2)$ [20, 34].
+This “Clifford + TT” paradigm underlies modern fault-tolerant quantum computation, where magic states enable non-Clifford operations through gate teleportation [7, 25].
 For higher-dimensional systems, the situation is less complete.
-Universality criteria for qudits of prime dimension have been established [Howard2012, Campbell2012], and the maximality of the Clifford group in these cases follows from classical results on finite linear groups [Nebe2001].
-However, composite dimensions with mixed prime factors have not been systematically studied, even as experimental platforms increasingly access non-binary levels [RingbauerEtAl2022, Low2025, Chicco_2023].
-Universality for arbitrary dimension $d$, not based on the Clifford group, was established through both constructive and algebraic frameworks.
-Previous work proved universality by decomposing elements of $\mathrm{SU}(d^{n})$ into two-level unitaries, which were further expressed through phase and controlled operations [LuoWang2014].
-Alternative constructions employed the generators of $\mathfrak{su}(d)$, where universal gate sets were generated via exponentiation of string operators, to produce $\mathrm{U}(d^{n})$ [VlasovJMP2002, VlasovSPIE2003].
-Main results. We classify the resources required for single-qudit universality by the prime factorization of $d$:
+Universality criteria for qudits of prime dimension have been established [23, 9], and the maximality of the Clifford group in these cases follows from classical results on finite linear groups [34].
+However, prime-power and composite dimensions have not been systematically studied, even as experimental platforms increasingly access non-binary levels [35, 31, 10].
+Universality for arbitrary dimension $d$, not based on the Clifford group, has been established through both constructive and algebraic frameworks.
+Previous work proved universality by decomposing elements of $\mathrm{SU}(d^{n})$ into two-level unitaries, which were further expressed through phase and controlled operations [32].
+Alternative constructions employed the generators of $\mathfrak{su}(d)$, where universal gate sets were generated via exponentiation of string operators, to produce $\mathrm{U}(d^{n})$ [39, 40].
+Main results.
+We provide a complete characterization of the resources required for single-qudit universality, based on the prime factorization of the local dimension $d$.
+The classification reflects a hierarchy of algebraic obstructions within the Clifford group:
   1. (I)
-    Prime dimensions $\bm{(d=p).}$ The Clifford group acts irreducibly on the Lie algebra and is a maximal finite subgroup of $\mathrm{PSU}(d)$.
-We prove maximality holds if and only if $d$ is prime: adding any non-Clifford gate forces universality.
-In this regime, every non-Clifford gate is automatically magic.
+    Prime dimensions $\bm{(d=p).}$ No obstruction.
+The Clifford group acts irreducibly on the Lie algebra and is a maximal finite subgroup of $\mathrm{SU}(d)$ modulo scalar phases.
+We prove that both maximality and irreducibility of the Clifford group hold if and only if $d$ is prime: adding any non-Clifford gate forces universality.
+In this case, every non-Clifford gate necessarily has magic.
   2. (II)
-    Prime power dimensions $\bm{(d=p^{m},\,m\geq 2).}$ The adjoint representation becomes reducible.
-The Lie algebra fragments into invariant subspaces indexed by divisibility.
-Universality requires specific diagonal non-Clifford gates—members of the $T_{s}$ family whose phases restore irreducibility.
-Not every non-Clifford gate suffices; the arithmetic obstruction must be overcome by deliberate design.
-Cases I and II share a common strategy: universality via diagonal non-Clifford gates (the $T_{s}$ family).
+    Prime-power dimensions $\bm{(d=p^{m},\,m\geq 2).}$ Reducibility obstruction.
+The adjoint representation becomes reducible, decomposing the Lie algebra into invariant subspaces indexed by divisibility.
+The Clifford group acts transitively within each subspace but cannot connect them.
+Universality requires a non-Clifford gate that restores irreducibility by coupling such subspaces.
+Two classes of gates accomplish this:
+    • Diagonal gates from the $T_{s}$ family with $s\nmid d$.
+    • Non-affine permutations, such as the transposition (0 1)(0\;1).
+    Not every non-Clifford gate suffices; one must choose a gate with the correct algebraic properties.
   3. (III)
-    Composite dimensions with coprime factors $\bm{(d=d_{1}\cdots d_{n}).}$ Universality emerges from intra-qudit gates alone.
-When $d$ has pairwise coprime factors, generalized CNOTs connecting subsystems construct diagonal gates via Bézout’s identity.
-These gates satisfy both density conditions: they restore irreducibility by mixing the local Clifford structures, and they certify infiniteness by lying within projective distance $1/2$ of the identity.
-No explicit magic injection is required.
-In contrast to prime dimensions, the failure of maximality allows for intermediate finite groups, meaning that a gate can be non-Clifford without being magic.
-Case III is our main contribution.
-In prime and prime-power dimensions (Case I and II), universality requires the explicit injection of diagonal non-Clifford gates ($T_{s}$).
-In contrast, coprime dimensions achieve universality using only generalized CNOTs.
-Although these are classical reversible permutations, they are non-Clifford resources.
-The arithmetic mismatch allows these permutation gates to generate a dense subgroup, making universality an intrinsic feature of the hardware that does not require explicit diagonal magic.
+    Composite dimensions with coprime factors $\bm{(d=d_{1}\cdots d_{k}).}$ Obstruction resolved by arithmetic.
+The Clifford group decomposes as a direct product $\mathcal{C}_{d}\cong\mathcal{C}_{d_{1}}\times\cdots\times\mathcal{C}_{d_{k}}$ of the local Clifford groups, and consequently fails to act irreducibly on the global Lie algebra (Proposition 8).
+However, generalized intra-qudit CNOTs connecting coprime factors automatically restore irreducibility and construct the necessary diagonal gates (Proposition 47).
+When $k=2$, a single CNOT coupling the two factors suffices; for $k>2$, multiple CNOTs connecting adjacent factors are required.
+These gates satisfy both density conditions: they restore irreducibility by coupling the local Clifford structures, and they certify infiniteness by lying within projective distance $1/2$ of the identity (Theorem 48).
+No explicit diagonal magic injection is required.
+Again, not every non-Clifford gate suffices (see conditions in Theorem 42).
+The trichotomy demonstrates that universality requirements depend strictly on how $d$ factorizes into prime powers.
+For prime dimensions, the Clifford group is maximally rigid—any non-Clifford gate suffices.
+For prime powers, reducibility demands gates that restore irreducibility.
+For coprime dimensions, intra-qudit CNOTs connecting the factors restore irreducibility and provide the necessary non-Clifford structure.
+In all three cases, the prime factorization determines both the obstruction to universality and the algebraic properties required to overcome it.
 Structure of the paper.
-Section II establishes the algebraic framework: the density criterion, the structure of the generalized Pauli and Clifford groups, and the decomposition of composite-dimensional systems via the Chinese Remainder Theorem.
-Section III develops the geometric criterion for infiniteness and its spectral characterization.
-Sections IV, V, and VI analyze the three regimes of the trichotomy in detail.
-Section VII discusses implications of our findings and open problems.
+Section II establishes our algebraic framework, including the density criterion based on the adjoint representation, the structure of the Clifford group via its symplectic quotient, and the factorization of composite systems.
+Section III introduces an infiniteness criterion: a subgroup with irreducible adjoint action is infinite if it contains a non-scalar element within projective distance $1/2$ of the identity.
+Sections IVVI systematically establish universality for each case of the trichotomy.
+Section VII concludes with implications and open questions.
 II Preliminaries
-II.1 Criteria for Universality
+II.1 Criterion for Universality
 The practical realization of quantum algorithms relies on the ability to approximate arbitrary unitary transformations using a finite set of elementary gates.
@@
 Once the single-qudit case has been completely characterized, extending this to universal quantum computation on an array of qudits requires only a single additional gate $V$ that breaks the preservation of the local tensor structure (an entangling resource).
-This reduction is formalized by the following seminal result:
-Theorem 2 ([Brylinski2002]).
+This reduction is formalized by the following result:
+Theorem 2 ([8]).
 Let $\mathcal{G}_{\mathrm{local}}\subset\mathrm{SU}(d)$ be a single-qudit universal set and $V\in\mathrm{SU}(d^{2})$ be a two-qudit gate.
@@
 Proving density directly via topological arguments is often hard.
-The fundamental difficulty lies in distinguishing whether a finitely generated subgroup is dense or merely trapped within a proper, infinite closed Lie subgroup of lower dimension.
+The main difficulty lies in distinguishing whether a finitely generated subgroup is dense or trapped within a proper, infinite closed Lie subgroup of lower dimension.
 Instead, we employ an algebraic criterion based on the rigidity of simple Lie algebras.
@@
 Since unitary conjugation preserves the Hilbert-Schmidt inner product, this action is unitary.
-A crucial property is that the irreducibility of this action is insensitive to complexification: the action on $\mathfrak{su}(d)$ is irreducible if and only if the complexified action on $\mathfrak{sl}(d,\mathbb{C})$ is irreducible.
+An important property is that the irreducibility of this action is insensitive to complexification: the action on $\mathfrak{su}(d)$ is irreducible if and only if the complexified action on $\mathfrak{sl}(d,\mathbb{C})$ is irreducible.
 We now establish the connection between the subgroup dimension and the representation.
@@
 ∎
-This lemma leads to one criteria for single-qudit universality.
+This lemma leads to the following criterion for single-qudit universality.
 Theorem 4.
@@
 $ZX=\omega XZ.$
-The Heisenberg group $H(d)$ (sometimes called the generalized Pauli group or just the Pauli group) consists of:
+The Heisenberg group $H(d)$ (sometimes called the generalized Pauli group or just the Pauli group [15, 12]) consists of:
 $H(d)=\begin{cases}\{\omega^{l}X^{a}Z^{b}:a,b,l\in\mathbb{Z}_{d}\}&\text{if }d\text{ is odd},\\ \{\zeta^{l}X^{a}Z^{b}:a,b\in\mathbb{Z}_{d},\,l\in\mathbb{Z}_{2d}\}&\text{if }d\text{ is even},\end{cases}$
@@
 This action corresponds to the special linear group over the ring $\mathbb{Z}/d\mathbb{Z}$.
-This structural rigidity is characterized by the short exact sequence [Tolar_2018]:
+This structural rigidity is characterized by the short exact sequence [37]:
 $1\longrightarrow\mathbb{Z}_{d}^{2}\longrightarrow C(d)\xrightarrow{\Psi}\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})\longrightarrow 1.$ (5)
@@
 $|\mathcal{C}_{d}|=|H(d)|\cdot|\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})|.$ (7)
-An important property for our analysis of composite systems is that this structure factorizes over coprime dimensions, allowing us to treat prime-power sectors independently.
+The finite Clifford group factorizes over coprime dimensions—a property previously identified by Looi and Griffiths [29] for stabilizer states—which allows us to reduce the problem to independent prime-power sectors.
 Proposition 8.
@@
 We show this maps generators to generators, inducing an isomorphism onto $\mathcal{C}_{d}$.
-Recall that by the Chinese Remainder Theorem (CRT), the map $\chi:\mathbb{Z}/d\mathbb{Z}\to\mathbb{Z}/{d_{1}}\mathbb{Z}\times\mathbb{Z}/{d_{2}}\mathbb{Z}$ given by $x\mapsto(x\bmod d_{1},x\bmod d_{2})$ is a ring isomorphism.
+Recall that by the Chinese Remainder Theorem, the map $\chi:\mathbb{Z}/d\mathbb{Z}\to\mathbb{Z}/{d_{1}}\mathbb{Z}\times\mathbb{Z}/{d_{2}}\mathbb{Z}$ given by $x\mapsto(x\bmod d_{1},x\bmod d_{2})$ is a ring isomorphism.
 Since the Clifford generators are defined entirely by arithmetic operations on the basis indices, the CRT implies their factorization:
@@
   2. 2.
-    Hadamard ($\bm{H}$): The Fourier transform kernel $\omega^{xy}$ factorizes because the inner product in $\mathbb{Z}_{d}$ decomposes into the sum of local inner products (modulo integer scaling), yielding $H_{d}\cong H_{d_{1}}\otimes H_{d_{2}}$.
+    Hadamard ($H$): The Fourier transform kernel $\omega^{xy}$ factorizes because the inner product in $\mathbb{Z}_{d}$ decomposes into the sum of local inner products (modulo integer scaling), yielding $H_{d}\cong H_{d_{1}}\otimes H_{d_{2}}$.
   3. 3.
-    Phase Gate ($\bm{P}$): The gate applies a phase $\omega^{Q(x)}$, where $Q(x)=x(x+\rho_{d})/2$ is a quadratic polynomial.
+    Phase Gate ($P$): The gate applies a phase $\omega^{Q(x)}$, where $Q(x)=x(x+\rho_{d})/2$ is a quadratic polynomial.
 Since $\chi$ is a ring homomorphism, it preserves polynomial evaluation: computing $Q(x)$ modulo $d$ is equivalent to computing pairs $(Q(x_{1})\bmod d_{1},Q(x_{2})\bmod d_{2})$.
@@
 Since the generators map bijectively and the groups have the same order, $\Phi$ is an isomorphism. ∎
-III A Geometric Criterion for Infiniteness
-Proving that a finitely generated matrix group is infinite is, in general, a non-trivial algebraic problem [detinko2013deciding].
+II.3 Permutations in the Clifford Group
+While the Clifford group acts mainly by creating superpositions via the Hadamard gate, it also contains a subgroup of classical reversible operations corresponding to permutations of the computational basis.
+Characterizing this subgroup is essential for identifying which classical gates constitute non-Clifford resources.
+Let $S_{d}$ be the symmetric group of degree $d$.
+For any permutation $\pi\in S_{d}$, we define the associated permutation matrix $P_{\pi}\in\mathrm{U}(d)$ by its action on the basis states: $P_{\pi}|x\rangle=|\pi(x)\rangle$.
+Proposition 9.
+A permutation matrix $P_{\pi}$ belongs to the Clifford group $\mathcal{C}_{d}$ if and only if the underlying permutation $\pi:\mathbb{Z}_{d}\to\mathbb{Z}_{d}$ is an invertible affine transformation over the ring $\mathbb{Z}_{d}$.
+That is, $\pi$ is of the form:
+$\pi(x)=ax+b\pmod{d}$, (8)
+where $a\in\mathbb{Z}_{d}^{\times}$ (a unit in $\mathbb{Z}_{d}$) and $b\in\mathbb{Z}_{d}$.
+Consequently, the subgroup of Clifford permutations is isomorphic to the affine group $\mathrm{Aff}(1,\mathbb{Z}_{d})$.
+Proof.
+$(\Rightarrow)$ Assume $P_{\pi}\in\mathcal{C}_{d}$. By definition, conjugation by $P_{\pi}$ must map the Heisenberg group $H(d)$ to itself.
+Consider the action on the Pauli-$Z$ operator, which is diagonal with entries $\omega^{x}$ where $\omega=e^{2\pi i/d}$. Conjugation permutes the diagonal entries:
+$P_{\pi}ZP_{\pi}^{\dagger}=\sum_{x\in\mathbb{Z}_{d}}\omega^{x}|\pi(x)\rangle\langle\pi(x)|=\sum_{y\in\mathbb{Z}_{d}}\omega^{\pi^{-1}(y)}|y\rangle\langle y|.$
+For the resulting diagonal operator to belong to $H(d)$, it must be of the form $\omega^{c_{0}}Z^{k}$ for some $k,c_{0}\in\mathbb{Z}_{d}$.
+Comparing the diagonal entries, we require $\omega^{\pi^{-1}(y)}=\omega^{c_{0}+ky}$ for all $y\in\mathbb{Z}_{d}$.
+This implies:
+$\pi^{-1}(y)\equiv ky+c_{0}\pmod{d}$.
+Since $\pi^{-1}$ is a bijection, the coefficient $k$ must be a unit in $\mathbb{Z}_{d}$; otherwise, the map $y\mapsto ky+c_{0}$ would fail to be injective.
+Thus, $\pi^{-1}$ is an affine map with invertible linear part, and consequently $\pi$ itself must be affine: $\pi(x)=ax+b$ with $a\in\mathbb{Z}_{d}^{\times}$.
+$(\Leftarrow)$ Conversely, assume $\pi(x)=ax+b$ with $a\in\mathbb{Z}_{d}^{\times}$. We verify the conjugation action on the generators $Z$ and $X$ of the Heisenberg group.
+For the Pauli-$Z$ operator, a direct calculation using $\pi^{-1}(y)=a^{-1}(y-b)$ gives:
+$P_{\pi}ZP_{\pi}^{\dagger}=\sum_{y\in\mathbb{Z}_{d}}\omega^{a^{-1}(y-b)}|y\rangle\langle y|=\omega^{-a^{-1}b}Z^{a^{-1}},$
+which belongs to $H(d)$.
+For the operator $X$, we compute:
+$P_{\pi}XP_{\pi}^{\dagger}|y\rangle$ $=P_{\pi}X|\pi^{-1}(y)\rangle=P_{\pi}|\pi^{-1}(y)+1\rangle$
+$=|\pi(\pi^{-1}(y)+1)\rangle.$
+Substituting $\pi^{-1}(y)=a^{-1}(y-b)$, the argument of $\pi$ becomes:
+$\pi\big(a^{-1}(y-b)+1\big)=a\cdot a^{-1}(y-b)+a+b=y+a.$
+Thus, $P_{\pi}XP_{\pi}^{\dagger}=X^{a}$. Since $a$ is a unit, $X^{a}$ generates the same cyclic subgroup as $X$ and belongs to $H(d)$.
+Since $P_{\pi}$ maps the generators of $H(d)$ to elements of $H(d)$, we conclude $P_{\pi}\in\mathcal{C}_{d}$. ∎
+Remark 10.
+The order of the affine group is $|\mathrm{Aff}(1,\mathbb{Z}_{d})|=\varphi(d)\cdot d$, where $\varphi$ denotes Euler’s totient function.
+For $d=2$, we have $|S_{2}|=2$ and $|\mathrm{Aff}(1,\mathbb{Z}_{2})|=1\cdot 2=2$.
+For $d=3$, we have $|S_{3}|=6$ and $|\mathrm{Aff}(1,\mathbb{Z}_{3})|=2\cdot 3=6$.
+In both cases, the groups coincide: every permutation is affine and thus a Clifford operation.
+However, for $d\geq 4$, the inequality $\varphi(d)\cdot d<d!$ implies the strict inclusion $\mathrm{Aff}(1,\mathbb{Z}_{d})\subsetneq S_{d}$.
+Consequently, non-affine permutations exist.
+For instance, in $d=4$, the transposition (0 1)(0\;1)—which swaps $|0\rangle$ and $|1\rangle$ while fixing $|2\rangle$ and $|3\rangle$—is a non-Clifford gate despite being a classical reversible operation.
+II.4 Diagonal Gates and the $T_{s}$-gate Family
+To characterize diagonal non-Clifford operators, for any dimension $d$, we introduce a generic family of diagonal gates.
+Definition 11.
+Let $\zeta:\mathbb{Z}_{d}\to\mathrm{U}(1)$ be a phase function. We define the associated diagonal gate $T_{\zeta}$ by its action on the computational basis:
+$T_{\zeta}|x\rangle=\zeta(x)|x\rangle,\quad\text{for }x\in\mathbb{Z}_{d}.$ (9)
+To determine whether a specific $T_{\zeta}$ belongs to the Clifford group, we examine its conjugation action on the operator $X$.
+A simple calculation shows that $T_{\zeta}XT_{\zeta}^{\dagger}=DX$, where $D$ is a diagonal operator encoding the relative phases.
+This motivates the introduction of the differential of the phase function.
+Definition 12.
+The differential (or 1-coboundary) of a function $\zeta:\mathbb{Z}_{d}\to\mathrm{U}(1)$ is the map $\delta_{\zeta}:\mathbb{Z}_{d}\times\mathbb{Z}_{d}\to\mathrm{U}(1)$ defined by:
+$\delta_{\zeta}(x,y)=\frac{\zeta(x+y)}{\zeta(x)\zeta(y)}.$ (10)
+Recall that a function $b:\mathbb{Z}_{d}\times\mathbb{Z}_{d}\to\mathrm{U}(1)$ is called a bicharacter if it is linear in both arguments.
+Lemma 13.
+Let $d$ be any integer. A diagonal gate $T_{\zeta}$ belongs to the Clifford group $\mathcal{C}_{d}$ if and only if its differential $\delta_{\zeta}$ is a bicharacter.
+Proof.
+The conjugation action on the shift operator $X^{y}$ is given by:
+$T_{\zeta}X^{y}T_{\zeta}^{\dagger}|k\rangle=\frac{\zeta(k)}{\zeta(k-y)}X^{y}|k\rangle.$
+For the result to be a generalized Pauli operator (of the form $Z^{a}X^{y}$), the coefficient function must be a character (linear in $k$).
+In terms of the differential, this condition requires $\delta_{\zeta}(k,-y)$ to be linear in $k$.
+Since this must hold for all $y$, $\delta_{\zeta}$ must be bilinear.
+Note that for the cyclic group $\mathbb{Z}_{d}$, any symmetric bicharacter takes the simple form $b(x,y)=\omega^{cxy}$, where $\omega=e^{2\pi i/d}$ is the primitive $d$-th root of unity and $c$ is an integer constant.
+Thus, the condition implies that $\zeta(x)$ must be a quadratic phase polynomial $\omega^{ax^{2}+bx}$.
+∎
+We specifically focus on the following $T_{s}$ family.
+Definition 14 (The $T_{s}$-gate Family).
+For any dimension $d\geq 2$ and any integer resolution $s\geq 2$, we define the diagonal gate $T_{s}$ acting on the computational basis as:
+$T_{s}=\sum_{k\in\mathbb{Z}_{d}}\exp\left(\frac{2\pi ik}{s}\right)|k\rangle\langle k|.$ (11)
+For instance, $T_{d}$ recovers the Pauli $Z$ gate (order $d$), and $T_{2d}$ corresponds to the generalized phase gate.
+To determine the “Cliffordness” of $T_{s}$, we analyze its conjugation action on $X$:
+$\operatorname{Ad}_{T_{s}}(X)=T_{s}XT_{s}^{\dagger}=\exp\left(\frac{2\pi i}{s}\right)X.$ (12)
+For $T_{s}$ to belong to the Clifford group $\mathcal{C}_{d}$, it must normalize the Heisenberg group.
+Since the result is proportional to $X$, this condition holds if and only if the phase $e^{2\pi i/s}I$ is in the Heisenberg group:
+  • Odd Dimensions: The phases are generated by $\omega=e^{2\pi i/d}$. Thus, $T_{s}\in\mathcal{C}_{d}\iff s\mid d$.
+  • Even Dimensions: The Clifford group generates phases of order $2d$. Thus, $T_{s}\in\mathcal{C}_{d}\iff s\mid 2d$.
+This leads to a unified criterion for when $T_{s}$ constitutes a non-Clifford resource.
+Proposition 15.
+The gate $T_{s}$ is a non-Clifford resource for $\mathrm{SU}(d)$ if and only if the order $s$ satisfies:
+$s\nmid K_{d},\quad\text{where }K_{d}=\begin{cases}d&\text{if }d\text{ is odd},\\ 2d&\text{if }d\text{ is even}.\end{cases}$ (13)
+III A Criterion for Infiniteness
+Proving that a finitely generated matrix group is infinite is, in general, a non-trivial algebraic problem [13].
 While the irreducibility of the adjoint action ensures that the group is not trapped in a proper Lie subgroup of lower dimension, it does not rule out the possibility that the group is finite.
 To address this, we introduce a topological condition based on the discrete nature of finite groups.
-Since we are interested in subgroups of $\mathrm{SU}(d)$, but physical operations are equivalent up to global phases, the appropriate topological setting is the projective unitary group $\text{PU}(d)=\mathrm{SU}(d)/Z(\mathrm{SU}(d))\cong\mathrm{U}(d)/\mathrm{U}(1)$.
+Since we are interested in subgroups of $\mathrm{SU}(d)$, but physical operations are equivalent up to global phases, the appropriate topological setting is the projective unitary group $\mathrm{PU}(d)=\mathrm{SU}(d)/Z(\mathrm{SU}(d))\cong\mathrm{U}(d)/\mathrm{U}(1)$.
 To capture the proximity to the identity in this quotient space, we employ a projective operator distance.
@@
 The projective distance of $U\in\mathrm{SU}(d)$ is defined as:
-$\mathrm{dist}_{\mathrm{proj}}(U,I)\coloneqq\min_{\phi\in[0,2\pi)}\|e^{i\phi}U-I\|,$ (8)
+$\mathrm{dist}_{\mathrm{proj}}(U,I)\coloneqq\min_{\phi\in[0,1)}\|e^{2\pi i\phi}U-I\|,$ (14)
 […]
-Note that the minimization is well-defined because the function $f(\phi)=\|e^{i\phi}U-I\|$ is continuous on the compact interval $0,2\pi]$.
-Also, while $U\in\mathrm{SU}(d)$, the optimal representative $e^{i\phi}U$ may lie in $\mathrm{U}(d)$.
+Note that the minimization is well-defined because the function $f(\phi)=\|e^{2\pi i\phi}U-I\|$ is continuous on the compact interval $[0,1]$.
+Also, while $U\in\mathrm{SU}(d)$, the optimal representative $e^{2\pi i\phi}U$ may lie in $\mathrm{U}(d)$.
 Theorem 17.
@@
 We now rely on a structural property established in the proof of Theorem 5.7 in Ref.
-Dixon1971: for any finite subgroup $G^{\prime}\subset\mathrm{GL}(d,\mathbb{C})$, the subgroup $A=\langle\{g\in G^{\prime}:\|g-I\|<1/2\}\rangle$ is a normal abelian subgroup.
+14: for any finite subgroup $G^{\prime}\subset\mathrm{GL}(d,\mathbb{C})$, the subgroup $A=\langle\{g\in G^{\prime}:\|g-I\|<1/2\}\rangle$ is a normal abelian subgroup.
 Since $h^{\prime}\in G^{\prime}$ satisfies $\|h^{\prime}-I\|<1/2$, it implies $h^{\prime}\in A$, so $A$ is non-trivial.
@@
 To adapt our infiniteness criterion to the projective setting, we identify the group of physical phases $U(1)$ with the quotient group $\mathbb{R}/\mathbb{Z}$ via the exponential map $\theta\mapsto e^{2\pi i\theta}$.
-To evaluate the projective distance efficiently, we first introduce a geometric measure of the angular concentration of the eigenvalues.
+To evaluate the projective distance efficiently, we first introduce a measure of the angular concentration of the eigenvalues.
 Definition 19.
@@
 We begin our analysis with the simplest case: dimensions $d=p$ where $p$ is a prime number.
-In this setting, the Clifford group exhibits a remarkable structural rigidity.
+In this setting, the Clifford group exhibits a strong structural rigidity.
 To understand this, we must first establish a general correspondence between the group’s action on the Lie algebra and the symplectic geometry of the discrete phase space.
-This framework will prove crucial not only for establishing universality in prime dimensions but also for diagnosing the obstructions that arise in prime-power dimensions (Sec.
+This framework will prove useful not only for establishing universality in prime dimensions but also for diagnosing the obstructions that arise in prime-power dimensions (Sec.
 V).
 IV.1 The Adjoint Action and $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ Orbits
-To explicitly decompose the adjoint representation of the Clifford group $C(d)$ on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$ into a direct sum of irreducible subrepresentations, we employ the theory of representations of group extensions developed by Clifford [clifford1937, Serre1977].
+To explicitly decompose the adjoint representation of the Clifford group $C(d)$ on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$ into a direct sum of irreducible subrepresentations, we employ the theory of representations of group extensions developed by Clifford [11, 36].
 The definition of the Clifford group implies that its adjoint action is monomial with respect to the generalized Pauli basis: elements of the group permute the basis vectors $\{V_{u}\}$ while applying phase factors.
@@
 IV.2 Irreducibility of the Adjoint Action
-The first condition of Theorem 4 is the irreducibility of the action on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$. We can now prove this characterization using the tools developed.
+The first condition of the density criterion (Theorem 4) is the irreducibility of the action on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$.
+As formalized by Graydon et al[21], a finite group forms a unitary 2-design if and only if its adjoint representation on the space of operators decomposes into exactly two irreducible components: the identity and the traceless subspace.
+While Graydon et al.
+demonstrated that this condition fails for composite dimensions (implying the Clifford group is not a 2-design in those cases), the following theorem establishes the positive counterpart: for prime dimensions, the adjoint action is strictly irreducible.
 Theorem 24.
@@
 The set of vectors with $\gcd(u,d)=k$ forms an orbit disjoint from the set of vectors with $\gcd(u,d)=1$.
-This implies the existence of at least a proper non zero invariant subspaces, rendering the representation reducible.
+This implies the existence of at least one proper nonzero invariant subspace, rendering the representation reducible.
 ∎
@@
 This property stems from the maximality of the Clifford group within the finite subgroups of the projective unitary group.
-While this fact is widely recognized in quantum information [Gottesman1999], a rigorous mathematical proof relies on deep results from finite group theory.
-For the general case of systems with $n$-qudits ($d$ prime), Nebe, Rains, and Sloane [Nebe2001] established maximality using sophisticated number-theoretic techniques.
+While this fact is well-known in quantum information [20], a rigorous mathematical proof relies on deep results from finite group theory.
+For the general case of systems with $n$-qudits ($d$ prime), Nebe, Rains, and Sloane [34] established maximality using number-theoretic techniques.
 Specifically, they utilized $l$-adic integrality and reduction modulo suitable rational primes to apply the classification of subgroups of finite general linear groups.
 However, as noted in Ref.
-[Nebe2001], for the single-qudit case (d=pd=p)—which is central to this work—the sufficiency of the primality condition follows from earlier work by Lindsey [Lindsey1970].
+[34], for the single-qudit case (d=pd=p)—which is central to this work—the sufficiency of the primality condition follows from earlier work by Lindsey [26].
 Unlike the general multi-qudit case, Lindsey’s proof relies primarily on character theory and the classification of linear groups of prime degree.
 Complementing this classical result, we demonstrate that primality is also a necessary condition, thereby establishing the following full characterization.
-For completeness, we provide a self-contained proof of the forward direction based on a Corollary in [Lindsey1970].
+For completeness, we provide a self-contained proof of the forward direction based on a Corollary in [26].
 Theorem 25.
@@
 We apply the following result of Lindsey:
-Theorem 19 ([Lindsey1970, Corollary, p. 58]).
+Theorem 26 ([26, Corollary, p. 58]).
 Let $G$ be a finite group with a faithful, primitive, irreducible representation of prime degree $p>5$.
@@
 If a finite extension $G\supsetneq\mathcal{C}_{p}$ existed, it would necessarily possess a Sylow $p$-subgroup of order at least $|\mathcal{C}_{p}|_{p}\geq p^{4}$ (accounting for the Heisenberg group and $SL(2,\mathbb{Z}/p\mathbb{Z})$ part).
-However, the classifications by Blichfeldt [Blichfeldt1917] and Brauer [Brauer1967] reveal that no primitive group—other than the normalizers of extraspecial groups—supports such a structure.
+However, the classifications by Blichfeldt [3] and Brauer [5] reveal that no primitive group—other than the normalizers of extraspecial groups—supports such a structure.
 For instance, for $p=5$, Brauer established that the only other primitive candidates involve central extensions of $A_{5},A_{6},$ or $\text{PSL}(2,11)$, whose Sylow $5$-subgroups have orders at most $25$, strictly less than the required order.
@@
 ∎
-IV.4 Universality and Diagonal Resources
+IV.4 Universality Resources in Prime Dimensions
 The combination of irreducibility (Theorem 24) and maximality (Theorem 25) yields an immediate criterion for universality, valid only in prime dimensions.
@@
 Let $p$ be a prime. Let $g\in\mathrm{SU}(p)$ be any unitary gate. If $g\notin\mathcal{C}_{p}$ (modulo scalars), then the set $\mathcal{C}_{p}\cup\{g\}$ is single-qudit universal.
-This corollary shifts the burden of finding universal sets to simply identifying a single non-Clifford gate. The most natural candidates are diagonal gates.
-Definition 21.
-Let $\zeta:\mathbb{Z}_{d}\to\mathrm{U}(1)$ be a phase function. We define the associated diagonal gate $T_{\zeta}$ by its action on the computational basis:
-$T_{\zeta}|x\rangle=\zeta(x)|x\rangle,\quad\text{for }x\in\mathbb{Z}_{d}.$ (13)
-To determine whether a specific $T_{\zeta}$ belongs to the Clifford group, we examine its conjugation action on the operator $X$.
-A simple calculation shows that $T_{\zeta}XT_{\zeta}^{\dagger}=DX$, where $D$ is a diagonal operator encoding the relative phases.
-This motivates the introduction of the differential of the phase function.
-Definition 22.
-The differential (or 1-coboundary) of a function $\zeta:\mathbb{Z}_{d}\to\mathrm{U}(1)$ is the map $\delta_{\zeta}:\mathbb{Z}_{d}\times\mathbb{Z}_{d}\to\mathrm{U}(1)$ defined by:
-$\delta_{\zeta}(x,y)=\frac{\zeta(x+y)}{\zeta(x)\zeta(y)}.$ (14)
-Recall that a function $b:\mathbb{Z}_{d}\times\mathbb{Z}_{d}\to\mathrm{U}(1)$ is called a bicharacter if it is linear in both arguments.
-Lemma 23.
-Let $d$ be any integer. A diagonal gate $T_{\zeta}$ belongs to the Clifford group $\mathcal{C}_{d}$ if and only if its differential $\delta_{\zeta}$ is a bicharacter.
+This corollary shifts the burden of finding universal sets to simply identifying a single non-Clifford gate.
+We can now identify specific resources based on the classifications established in the Preliminaries.
+First, regarding diagonal resources:
+Corollary 28.
+Let $p$ be a prime number. The set $\mathcal{C}_{p}\cup\{T_{s}\}$ is single-qudit universal if and only if $s\nmid K_{p}$ (as defined in Proposition 15).
+Second, we consider permutation resources.
+Since the Clifford group contains only affine permutations (Proposition 9), any permutation that violates the affine condition $x\mapsto ax+b$ constitutes a universal resource.
+Corollary 29.
+Let $p\geq 5$ be a prime.
+Any non-affine permutation $\pi\in S_{p}$ yields a universal gate set $\mathcal{C}_{p}\cup\{P_{\pi}\}$.
+In particular, the simple transposition $\tau=(0\;1)$, which swaps the first two basis states and fixes the rest, is universal for all $p\geq 5$.
 Proof.
-The conjugation action on the shift operator $X^{y}$ is given by:
-$T_{\zeta}X^{y}T_{\zeta}^{\dagger}|k\rangle=\frac{\zeta(k)}{\zeta(k-y)}X^{y}|k\rangle.$
-For the result to be a generalized Pauli operator (of the form $Z^{a}X^{y}$), the coefficient function must be a character (linear in $k$).
-In terms of the differential, this condition requires $\delta_{\zeta}(k,-y)$ to be linear in $k$.
-Since this must hold for all $y$, $\delta_{\zeta}$ must be bilinear.
-Note that for the cyclic group $\mathbb{Z}_{d}$, any symmetric bicharacter takes the simple form $b(x,y)=\omega^{cxy}$, where $\omega=e^{2\pi i/d}$ is the primitive $d$-th root of unity and $c$ is an integer constant.
-Thus, the condition implies that $\zeta(x)$ must be a quadratic phase polynomial $\omega^{ax^{2}+bx}$.
+By Corollary 27, it suffices to show $P_{\tau}\notin\mathcal{C}_{p}$, which holds if $\tau$ is not affine.
+Suppose $\tau(x)=ax+b$.
+Evaluated at $x=0$, $\tau(0)=1\implies b=1$.
+Evaluated at $x=1$, $\tau(1)=0\implies a(1)+1=0\implies a=-1$.
+Evaluated at $x=2$ (since $p\geq 3$, the element 2 exists), $\tau(2)=2$.
+However, the affine formula predicts $\tau(2)=(-1)(2)+1=-1$.
+Thus, equality holds only if $2\equiv-1\pmod{p}$, i.e., $3\equiv 0\pmod{p}$.
+This is possible only for $p=3$.
+For all primes $p\geq 5$, the transposition is non-affine and therefore universal.
 ∎
-This lemma provides a constructive method to generate universal gate sets.
-For a prime dimension $p$, any diagonal gate whose differential is not a bicharacter (corresponding to a non-quadratic phase) is universal.
-IV.5 The $T_{s}$-gate Family
-The criteria for universality derived in the previous sections—maximality for prime dimensions and irreducibility via mixing for composite dimensions—share a common requirement: the availability of a unitary that lies strictly outside the local Clifford group.
-To characterize such operators, for any dimension $d$, we introduce a generic family of diagonal gates.
-Definition 24 (The $T_{s}$-gate Family).
-For any dimension $d\geq 2$ and any integer resolution $s\geq 2$, we define the diagonal gate $T_{s}$ acting on the computational basis as:
-$T_{s}=\sum_{k\in\mathbb{Z}_{d}}\exp\left(\frac{2\pi ik}{s}\right)|k\rangle\langle k|.$ (15)
-This family generalizes the standard Clifford hierarchy. For instance, $T_{d}$ recovers the Pauli $Z$ gate (order $d$), and $T_{2d}$ corresponds to the generalized phase gate.
-To determine the “Cliffordness” of $T_{s}$ for an arbitrary dimension, we analyze its conjugation action on the Pauli shift operator $X$. A direct calculation yields:
-$\operatorname{Ad}_{T_{s}}(X)=T_{s}XT_{s}^{\dagger}=\exp\left(\frac{2\pi i}{s}\right)X.$ (16)
-For $T_{s}$ to belong to the Clifford group $\mathcal{C}_{d}$, it must normalize the Heisenberg group.
-Since the result is proportional to the operator $X$, this condition holds if and only if the $e^{2\pi i/s}I$ is in the Heisenberg group:
-  • Odd Dimensions: The phases are generated by the primitive $d$-th root of unity $\omega=e^{2\pi i/d}$. Thus, $T_{s}\in\mathcal{C}_{d}\iff s\mid d$.
-  • Even Dimensions: The Clifford group generates phases of order $2d$.
-Specifically for qubits ($d=2$), the group includes the order-4 phase $i$.
-Thus, generally, $T_{s}\in\mathcal{C}_{d}\iff s\mid 2d$.
-The previous discussion leads to a unified criterion for when $T_{s}$ constitutes a non-Clifford resource.
-Proposition 25.
-The gate $T_{s}$ is a non-Clifford resource for $\mathrm{SU}(d)$ if and only if the order $s$ satisfies:
-$s\nmid K_{d},\quad\text{where }K_{d}=\begin{cases}d&\text{if }d\text{ is odd},\\ 2d&\text{if }d\text{ is even}.\end{cases}$ (17)
-Corollary 26.
-Let $p$ be a prime number. The set $\mathcal{C}_{p}\cup\{T_{s}\}$ is single-qudit universal if and only if $s\nmid K_{p}$.
-V Case II: Prime Power Dimensions
-While prime dimensions provide a rigid structure where the Clifford group is maximal and irreducible, prime power dimensions $d=p^{m}$ (with $m\geq 2$) introduce a rich internal structure derived from the arithmetic of the ring $\mathbb{Z}/{p^{m}}\mathbb{Z}$.
-In this regime, the two primary pillars of universality no longer hold: the adjoint action becomes reducible, and the group loses its maximality.
+V Case II: Prime-power Dimensions
+While prime dimensions provide a rigid structure where the Clifford group is projectively finite maximal and irreducible, prime-power dimensions $d=p^{m}$ (with $m\geq 2$) introduce a structure derived from the arithmetic of the ring $\mathbb{Z}/{p^{m}}\mathbb{Z}$.
+In this case, the two primary pillars of universality no longer hold: the adjoint action becomes reducible, and the group loses its maximality.
+To achieve universality, one must add a gate that restores irreducibility while ensuring infiniteness.
+We show that two classes of non-Clifford gates accomplish this: diagonal gates from the $T_{s}$ family (Sec.
+V.3) and non-affine permutations such as transpositions (Sec.
+V.4).
 V.1 Reducibility and Orbit Decomposition
@@
 IV, we established that the irreducibility of the Clifford group’s action is equivalent to the transitivity of the $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ group on the Pauli indices (Lemma 22).
-While the prime case ($d=p$) guarantees transitivity, the prime power case ($d=p^{m}$, $m\geq 2$) introduces structural modifications due to the existence of non-trivial ideals in the ring $\mathbb{Z}/{p^{m}}\mathbb{Z}$.
+While the prime case ($d=p$) guarantees transitivity, the prime-power case ($d=p^{m}$, $m\geq 2$) introduces structural modifications due to the existence of non-trivial ideals in the ring $\mathbb{Z}/{p^{m}}\mathbb{Z}$.
 Applying the machinery established in Lemma 22 and Lemma 23, we identify that the action of $\mathrm{SL}(2,\mathbb{Z}_{p^{m}})$ preserves the valuation of the index vector.
@@
 Theorem 30.
-For a prime power dimension $d=p^{m}$ ($m\geq 2$), the adjoint representation of the Clifford group on $\mathfrak{sl}(d,\mathbb{C})$ is reducible.
+For a prime-power dimension $d=p^{m}$ ($m\geq 2$), the adjoint representation of the Clifford group on $\mathfrak{sl}(d,\mathbb{C})$ is reducible.
 The algebra decomposes into a direct sum of $m$ irreducible orthogonal invariant subspaces:
@@
 $\mathcal{W}_{k}=\operatorname{span}\left\{P_{(u,v)}:\gcd(u,v,p^{m})=p^{k}\right\}.$ (20)
-This result fundamentally shifts the criteria for universality compared to the prime case.
+This result changes the criterion for universality compared to the prime case.
 Since the Lie algebra is fragmented into irreducible subspaces $\mathcal{W}_{k}$ that the Clifford group cannot connect, a supplementary gate must do more than simply generate infinite order; it must actively “mix” these subspaces, bridging irreducibility.
@@
 Consider the diagonal unitary gate $T_{\zeta}$ associated with a phase function $\zeta:\mathbb{Z}_{d}\to\mathrm{U}(1)$, as defined in Sec. IV.4. The adjoint action on an operator $X^{a}$ is given by:
-$\operatorname{Ad}_{T_{\zeta}}(X^{a})=\zeta(a)X^{a}D_{a},$ (20)
-where $D_{a}$ is the diagonal operator with entries $(D_{a})_{xx}=\delta_{\zeta}(a,x)$. Here, $\delta_{\zeta}$ is the differential previously introduced in Eq. (14).
+$\operatorname{Ad}_{T_{\zeta}}(X^{a})=X^{a}D_{a},$ (21)
+where $D_{a}$ is the diagonal operator with entries
+$(D_{a})_{xx}=\frac{\zeta(a+x)}{\zeta(x)}.$ (22)
+Equivalently, $(D_{a})_{xx}=\zeta(a)\,\delta_{\zeta}(a,x)$, where $\delta_{\zeta}$ is the 1-coboundary previously introduced in Eq. (10).
 To identify which subspaces $\mathcal{W}_{k}$ are populated by this operation, we expand the diagonal operator $D_{a}$ in the Pauli-$Z$ basis using the Discrete Fourier Transform (DFT).
@@
 Substituting the inverse DFT of the differential row $x\mapsto\delta_{\zeta}(a,x)$, the adjoint action expands in terms of the generalized Pauli basis $V_{(a,b)}=X^{a}Z^{b}$ as follows:
-$\operatorname{Ad}_{T_{\zeta}}(X^{a})=\frac{\zeta(a)}{\sqrt{d}}\sum_{b\in\mathbb{Z}_{d}}\widehat{(\delta_{\zeta,a})}(b)V_{(a,b)}.$ (22)
+$\operatorname{Ad}_{T_{\zeta}}(X^{a})=\frac{1}{\sqrt{d}}\sum_{b\in\mathbb{Z}_{d}}\widehat{(\delta_{\zeta,a})}(b)V_{(a,b)}.$ (24)
 This expansion reveals the selection rules of the gate.
@@
 (24) and the Orbit-Mixing Condition (Def.
-28), the expansion of this vector contains a term $\beta V_{(p^{k},b)}$ with $\beta\neq 0$ and $b\in\mathbb{Z}_{d}^{\times}$.
+31), the expansion of this vector contains a term $\beta V_{(p^{k},b)}$ with $\beta\neq 0$ and $b\in\mathbb{Z}_{d}^{\times}$.
 Since $b$ is a unit, $\gcd(p^{k},b,p^{m})=1$, which implies $V_{(p^{k},b)}\in\mathcal{W}_{0}$.
@@
 V.3 Universality via the $T_{s}$-gate Family
-We now apply our general machinery—irreducibility via orbit-mixing gates and infiniteness via the geometric criterion—to the generic family of diagonal gates $T_{s}$.
+We now apply our general machinery—irreducibility via orbit-mixing gates and infiniteness via the infiniteness criterion—to the generic family of diagonal gates $T_{s}$.
 Recall that $T_{s}$ introduces phases proportional to the roots of unity of order $s$, see Eq.
 (11).
+Note that $T_{s}$ is a unitary diagonal operator (hence $T_{s}\in\mathrm{U}(d)$), and it may fail to have determinant one.
+Since our density and infiniteness criteria are projective (and the adjoint action is insensitive to global phases), we may freely replace $T_{s}$ by any scalar multiple in $\mathrm{SU}(d)$ without affecting the arguments.
 Theorem 33.
@@
     The group $G_{s}$ is single-qudit universal (dense in $\mathrm{SU}(d)$) if $s\nmid d$ and $s$ satisfies the bound:
-    $s>\frac{\pi(d-1)}{2\arcsin(1/4)}.$ (24)
+    $s>\pi(d-1)\big/(2\,\arcsin(1/4)).$ (26)
 Proof.
@@
 If this bound holds, the group is infinite. Combined with the irreducibility from Item 1, the group is dense in $\mathrm{SU}(d)$. ∎
+V.4 Permutations as Non-Clifford Resources
+While diagonal gates of the $T_{s}$ family provide a method to achieve universality, permutation gates offer an alternative class of non-Clifford resources.
+In this subsection, we analyze when non-affine permutations can supplement the Clifford group to generate a dense subgroup of $\mathrm{SU}(d)$.
+V.4.1 Induced Diagonal Operators
+Conjugation by a permutation matrix transforms the Pauli-$Z$ operator into a diagonal operator that encodes the permutation structure.
+Explicitly, the adjoint action yields the diagonal operator $D_{\pi}$ given by:
+$P_{\pi}ZP_{\pi}^{\dagger}=D_{\pi},\quad\text{where}\quad(D_{\pi})_{yy}=\omega^{\pi^{-1}(y)}.$ (30)
+Since $D_{\pi}$ and $Z$ are related by unitary conjugation, they share the same spectrum: the complete set of $d$-th roots of unity.
+Consequently, $D_{\pi}$ has maximal spectral span $\ell(D_{\pi})=(d-1)/d$ and cannot directly satisfy the infiniteness criterion of Theorem 17.
+To extract a useful resource, we consider the difference operator that isolates the deviation from the identity permutation.
+Definition 34.
+For a permutation $\pi\in S_{d}$, the difference operator is:
+$h_{\pi}=D_{\pi}Z^{\dagger}=P_{\pi}ZP_{\pi}^{\dagger}Z^{\dagger},$ (31)
+with diagonal entries $(h_{\pi})_{yy}=\omega^{\pi^{-1}(y)-y}$.
+The difference operator measures how much $\pi$ displaces each index. For an affine permutation $\pi(x)=ax+b$, we have $\pi^{-1}(y)=a^{-1}(y-b)$, so the exponent becomes:
+$\pi^{-1}(y)-y=(a^{-1}-1)y-a^{-1}b,$
+which is linear in $y$. Thus, $h_{\pi}\in H(d)$ for affine permutations. For non-affine permutations, $h_{\pi}$ exhibits a non-linear phase structure.
+V.4.2 Transpositions and the Infiniteness Criterion
+We analyze the simplest non-affine permutations: transpositions.
+Proposition 35.
+Let $\tau=(j\;k)$ be the transposition swapping indices $j$ and $k$ with $j<k$. The difference operator is:
+$h_{\tau}=\mathrm{diag}(1,\ldots,1,\omega^{k-j},1,\ldots,1,\omega^{-(k-j)},1,\ldots,1),$ (32)
+where the entries $\omega^{k-j}$ and $\omega^{-(k-j)}$ occur at positions $j$ and $k$, respectively. The spectral span is:
+$\ell(h_{\tau})=\frac{2(k-j)}{d}.$ (33)
+Proof.
+Since transpositions are self-inverse, $\tau^{-1}=\tau$. The diagonal entries of $h_{\tau}$ are $(h_{\tau})_{yy}=\omega^{\tau(y)-y}$:
+  • For $y=j$: $\tau(j)=k$, so $(h_{\tau})_{jj}=\omega^{k-j}$.
+  • For $y=k$: $\tau(k)=j$, so $(h_{\tau})_{kk}=\omega^{j-k}=\omega^{-(k-j)}$.
+  • For $y\notin\{j,k\}$: $\tau(y)=y$, so $(h_{\tau})_{yy}=1$.
+The eigenvalues are $\{1,\omega^{k-j},\omega^{-(k-j)}\}$ with multiplicities d2d-2, 11, and $1$, respectively.
+These span an arc from angle $2\pi(k-j)/d$ to $2\pi(k-j)/d$, giving spectral span $\ell(h_{\tau})=2(k-j)/d$.
+∎
+Corollary 36.
+The transposition $\tau=(j\;k)$ generates an infinite group $\langle\mathcal{C}_{d},P_{\tau}\rangle$ provided:
+$k-j<\frac{d\cdot\arcsin(1/4)}{\pi}\approx 0.0805\cdot d.$ (34)
+In particular, the adjacent transposition $\tau=(0\;1)$ satisfies this bound for all $d\geq 13$.
+Proof.
+By Lemma 20, the projective distance is:
+$\mathrm{dist}_{\mathrm{proj}}(h_{\tau},I)=2\sin\left(\frac{\pi\ell(h_{\tau})}{2}\right)=2\sin\left(\frac{\pi(k-j)}{d}\right).$
+The infiniteness criterion (Theorem 17) requires this distance to be less than $1/2$:
+$2\sin\left(\frac{\pi(k-j)}{d}\right)<\frac{1}{2}\implies\sin\left(\frac{\pi(k-j)}{d}\right)<\frac{1}{4}.$
+For the adjacent transposition with kj=1k-j=1, this becomes $\sin(\pi/d)<1/4$, which holds when $d>\pi/\arcsin(1/4)\approx 12.43$. ∎
+V.4.3 Irreducibility Analysis for Transpositions
+We now verify that the difference operator $h_{\tau}$ induced by the adjacent transposition satisfies the orbit-mixing condition of Definition 31.
+Let $d=p^{m}$ with $m\geq 2$, and let $\tau=(0\;1)$.
+The phase function $\zeta:\mathbb{Z}_{d}\to\mathrm{U}(1)$ of the difference operator $h_{\tau}$ is given by $\zeta(0)=\omega$, $\zeta(1)=\omega^{-1}$, and $\zeta(y)=1$ for $y\geq 2$.
+For any proper divisor $u=p^{k}$ with $1\leq k\leq m-1$, a direct calculation of the differential row $f_{u}(y)=\zeta(u+y)/\zeta(y)$ yields:
+$f_{u}(y)=\begin{cases}\omega^{-1}&y\in\{0,\,d-u+1\},\\ \omega&y\in\{1,\,d-u\},\\ 1&\text{otherwise},\end{cases}$ (35)
+where indices are taken modulo $d$. Using this explicit form, we determine the spectral properties of the gate.
+Proposition 37.
+For $n\not\equiv 0\pmod{d}$, the Discrete Fourier Transform of the differential row $f_{u}$ is:
+$\sqrt{d}\,\hat{f}_{u}(n)$ $=(\omega^{-1}-1)(1+\omega^{n(u-1)})$
+$\quad+(\omega-1)(\omega^{-n}+\omega^{nu}).$ (36)
+Proof.
+The Fourier transform is $\sqrt{d}\,\hat{f}_{u}(n)=\sum_{y=0}^{d-1}f_{u}(y)\,\omega^{-ny}$. We decompose this sum by separating the deviation from the constant function:
+$\sqrt{d}\,\hat{f}_{u}(n)=\sum_{y=0}^{d-1}\omega^{-ny}+\sum_{y\in S}(f_{u}(y)-1)\omega^{-ny},$
+where $S=\{0,1,d-u,d-u+1\}$ is the support of the deviation.
+For $n\not\equiv 0\pmod{d}$, the first sum vanishes: $\sum_{y=0}^{d-1}\omega^{-ny}=0$.
+The second sum evaluates to:
+$(\omega^{-1}-1)\cdot\omega^{0}+(\omega-1)\cdot\omega^{-n}$
+$+(\omega-1)\cdot\omega^{-n(d-u)}+(\omega^{-1}-1)\cdot\omega^{-n(d-u+1)}.$
+Using $\omega^{d}=1$, we simplify:
+$\omega^{-n(d-u)}$ $=\omega^{-nd+nu}=\omega^{nu},$
+$\omega^{-n(d-u+1)}$ $=\omega^{-nd+nu-n}=\omega^{n(u-1)}.$
+Substituting and grouping terms yields Eq. (37). ∎
+Proposition 38.
+Let $d=p^{m}$ with $m\geq 2$. The difference operator $h_{\tau}$ for $\tau=(0\;1)$ is an orbit-mixing gate in the sense of Definition 31.
+Proof.
+We must show that for each proper divisor $u=p^{k}$ ($1\leq k\leq m-1$), there exists at least one unit $n\in\mathbb{Z}_{d}^{\times}$ such that $\hat{f}_{u}(n)\neq 0$.
+We prove that $n=1$ always works.
+Suppose, for contradiction, that $\hat{f}_{u}(1)=0$. From Eq. (37) with $n=1$:
+$(\omega^{-1}-1)(1+\omega^{u-1})+(\omega-1)(\omega^{-1}+\omega^{u})=0.$
+Using the identity $\omega^{-1}-1=-\omega^{-1}(\omega-1)$ and dividing by $(\omega-1)\neq 0$:
+$\omega^{-1}(1+\omega^{u-1})+(\omega^{-1}+\omega^{u})=0.$
+Expanding and simplifying:
+$\omega^{-1}-\omega^{u-2}+\omega^{-1}+\omega^{u}=0,$
+which reduces to $\omega^{u}=\omega^{u-2}$, i.e., $\omega^{2}=1$.
+Since $\omega=e^{2\pi i/d}$, this requires $d\mid 2$.
+For any prime power $d=p^{m}\geq 4$ with $m\geq 2$, this condition fails.
+Therefore, $\hat{f}_{u}(1)\neq 0$ for every proper divisor $u$, confirming the orbit-mixing condition.
+∎
+V.4.4 Universality Theorem for Transpositions
+Combining the results on infiniteness and irreducibility, we obtain:
+Theorem 39.
+Let $d=p^{m}$ with $m\geq 2$. The group $G=\langle\mathcal{C}_{d},P_{\tau}\rangle$, where $\tau=(0\;1)$ is the adjacent transposition, is dense in $\mathrm{SU}(d)$.
+Proof.
+By Theorem 4, it suffices to verify irreducibility of the adjoint representation and infiniteness $G$.
+Irreducibility: The difference operator $h_{\tau}=P_{\tau}ZP_{\tau}^{\dagger}Z^{\dagger}$ belongs to $G$.
+By Proposition 38, $h_{\tau}$ is an orbit-mixing gate.
+Theorem 32 then implies that $G$ acts irreducibly on $\mathfrak{sl}(d,\mathbb{C})$.
+Infiniteness: For $d\geq 13$, Corollary 36 establishes that the projective distance satisfies $\mathrm{dist}_{\mathrm{proj}}(h_{\tau},I)<1/2$, and Theorem 17 implies that $G$ is infinite.
+For the remaining prime-power cases $d\in\{4,8,9\}$, infiniteness has been verified by explicit computation using the computer algebra system Magma [4].
+The following code confirms that $\langle\mathcal{C}_{d},P_{\tau}\rangle$ is infinite in each case:
+1for d in [4, 8, 9] do
+2 ord := LCM(d, 2*d);
+3 F := CyclotomicField(ord);
+4 z := F.1; omega := z^(ord div d);
+5 zeta := z^(ord div (2*d));
+6 H := (1/F!Isqrt(d)) *
+7 Matrix(F,d,d,[omega^(j*k): j,k in [0..d-1]]);
+8 P := DiagonalMatrix(F,
+9 [zeta^(j*(j + d mod 2)): j in [0..d-1]]);
+10 X := Matrix(F,d,d,[<i,(i mod d)+1,1>: i in [1..d]]);
+11 tau := IdentityMatrix(F, d);
+12 tau[1,1]:=0; tau[1,2]:=1; tau[2,1]:=1; tau[2,2]:=0;
+13 G := sub<GL(d, F) | H, P, X, tau>;
+[…]
+15end for;
+16
+17// Output: d=4: Infinite=true
+18// d=8: Infinite=true
+19// d=9: Infinite=true
+The two conditions of Theorem 4 are satisfied for all prime powers $d=p^{m}$ with $m\geq 2$, completing the proof. ∎
 VI Case III: Composite Dimensions
-We now turn to the final regime: composite dimensions $d=d_{1}d_{2}\cdots d_{n}$ where the factors are pairwise coprime.
+We now turn to the final case: composite dimensions $d=d_{1}d_{2}\cdots d_{n}$ where the factors are pairwise coprime.
 Unlike prime or prime-power dimensions, which typically require the direct injection of diagonal non-Clifford gates (such as the qubit $T$-gate or the qudit $T_{s}$), composite dimensions offer a path to universality driven by “structural mixing.”
-In this regime, the necessary resources are not diagonal non-Clifford gates, but rather standard entangling gates (generalized CNOTs) that couple different coprime factors: intra-qudit gates.
+In this case, the necessary resources are not diagonal non-Clifford gates, but rather standard entangling gates (generalized CNOTs) that couple different coprime factors: intra-qudit gates.
 While these coupling gates are external to the local Clifford product, they are “natural” resources in any register-based architecture: they are permutation matrices that assist entanglement generation.
-The central result of this Section is the demonstration that the arithmetic interplay between coprime dimensions, coupled by intra-qudit gates, spontaneously generates the non-Clifford diagonal resources required for universality.
+The central result of this Section is the demonstration that the arithmetic interplay between coprime dimensions, coupled by intra-qudit gates, automatically generates the non-Clifford diagonal resources required for universality.
 VI.1 Irreducibility via Tensor Mixing
@@
     If $k=l$, $V\notin\mathrm{SU}(k)\otimes\mathrm{SU}(k)$ and $V\notin(\mathrm{SU}(k)\otimes\mathrm{SU}(k))\cdot S$.
-This result generalizes the seminal theorem of Brylinski and Brylinski [Brylinski2002], which established that adjoining any entangling gate to the local unitary group $\mathrm{SU}(d)\otimes\mathrm{SU}(d)$ yields a universal set for two-qudit quantum computation.
+This result generalizes a theorem by Brylinski and Brylinski [8], which established that adjoining any entangling gate to the local unitary group $\mathrm{SU}(d)\otimes\mathrm{SU}(d)$ yields a universal set for two-qudit quantum computation.
 VI.2 Intra-qudit Gates and Induced Magic
@@
 Theorem 48.
-Let $d=d_{1}d_{2}\cdots d_{n}$ be a proper composite dimension decomposed into pairwise coprime prime power factors (i.e., $d_{k}=p_{k}^{m_{k}}$ with $\gcd(d_{i},d_{j})=1$ for $i\neq j$).
+Let $d=d_{1}d_{2}\cdots d_{n}$ be a proper composite dimension decomposed into pairwise coprime prime-power factors (i.e., $d_{k}=p_{k}^{m_{k}}$ with $\gcd(d_{i},d_{j})=1$ for $i\neq j$).
 The group generated by the single-qudit Clifford group and the intra-qudit gates; that is, the group
@@
 By Theorem 33, the group is dense on factor $k$ if $S=d_{k}d_{j}$ satisfies the bound:
-    $S>\frac{\pi(d_{k}-1)}{2\arcsin(1/4)}.$
+    $S>\pi(d_{k}-1)\big/(2\,\arcsin(1/4)).$
     Consider the worst-case scenario with the smallest possible coprime proper powers: $d_{k}=4$ and $d_{j}=9$.
@@
 Since $36>18.66$, the condition is strictly satisfied.
-Any other pair of proper powers yields a larger product, satisfying the bound even more robustly.
+Any other pair of proper powers yields a larger product, satisfying the bound strictly.
 Therefore, the group is dense in the subsystem $k$.
@@
 A consequence of Theorem 48 is that for composite systems with coprime dimensions, universality is achieved using only standard architectural primitives.
-In the qubit regime ($d=2$), the group generated by local Clifford gates and CNOTs remains finite (the global Clifford group), requiring the injection of a dedicated magic resource (such as a $T$-gate) to achieve density.
+In the qubit case ($d=2$), the group generated by local Clifford gates and CNOTs remains finite (the global Clifford group), requiring the injection of a dedicated magic resource (such as a $T$-gate) to achieve density.
 In contrast, for coprime composite systems, the inclusion of the intra-qudit gate CNCN—typically considered a standard permutation—is sufficient to break the finite group structure.
@@
 Corollary 49.
-Let $d=d_{1}d_{2}\cdots d_{n}$ be a proper composite dimension with pairwise coprime prime power factors.
+Let $d=d_{1}d_{2}\cdots d_{n}$ be a proper composite dimension with pairwise coprime prime-power factors.
 The following finite set of gates constitutes a universal set for quantum computation on arrays of registers of dimension $d$:
@@
 VII Discussion and Outlook
-The results of this work show that the resources required to supplement the single-qudit Clifford group for universal quantum computation depend fundamentally on the arithmetic structure of the Hilbert space dimension.
-This dependence is not merely quantitative—involving more or fewer extra gates—but also qualitative: the type of resource changes across the trichotomy.
-This establishes dimensionality as a computational resource: the factorization properties of the dimension influence not only the cost but the nature of universality.
-Our perspective reframes the universality question.
-Rather than asking which gates generate a dense subgroup of $\mathrm{SU}(d)$, we ask which operations compensate for structural limitations of the Clifford group.
-Gates that are merely non-Clifford never suffice unless $d=p$, where non-Cliffordness implies magic.
-Only those that have genuinely “magic” guarantee single-qudit universality.
-In composite dimensions, these notions diverge: adding a non-Clifford gate can enlarge the group without making it infinite, leaving universality unattained.
-Achieving universality, therefore, requires gates that both escape the Clifford group and ensure infiniteness.
-The failure of density may arise from reducibility of the adjoint action, finiteness of the group, or both—and the prime factorization of the dimension dictates which condition applies.
-The classification we have presented reveals an unexpected connection between number theory and quantum computation.
-The prime factorization of the Hilbert space dimension determines whether universality requires explicit magic injection or emerges from structure.
-This insight highlights dimensionality as a design lever: by choosing coprime or prime-power architectures, one can trade off magic resources against control complexity.
-This connection is not merely formal; it may have concrete implications for the design of quantum processors.
-The broader lesson is that quantum resources are context-dependent.
-Entanglement, magic, and coherence are necessary for quantum advantage, but which resources are “free” and which must be engineered depends on the computational framework.
-Dimensionality is context-dependent as well.
-Its arithmetic structure governs whether magic is intrinsic or must be injected.
-In homogeneous qubit architectures, entangling gates are free but magic is costly.
-In coprime architectures, both are free—the price to pay lies in the complexity of controlling heterogeneous subsystems.
-Understanding these trade-offs across different physical platforms is essential for identifying the most efficient routes to practical quantum computation.
-Several limitations and open questions remain.
-For instance, multi-qudit universality, i.e.
-density in $\mathrm{SU}(d^{n})$, additionally requires an entangling gate satisfying the Brylinski criterion [Brylinski2002].
-For coprime architectures, intra-qudit CNOT gates can serve as both magic and entangling resources, though compilation overhead in heterogeneous registers remains an open question.
-Also, realizing coprime dimensions physically poses challenges: high-fidelity gates between subsystems of different dimensions, such as a generalized intra-qudit CNOT between a qutrit and a ququart, demand precise control of a twelve-dimensional Hilbert space.
-On the other hand, resource-theoretic measures of magic-such as stabilizer nullity [Beverland_2020], mana [Veitch2014], robustness [Howard2017], and stabilizer extent [Bravyi_2019] must be understood when magic emerges from coupling rather than injection.
-Beyond qudits, whether analogues of our trichotomy exist in continuous-variable systems, which have their own Clifford group and magic resources [Bartlett2002], or in hybrid architectures, is an intriguing question.
-Additionally, the complexity-theoretic implications of universality in the composite case remain unclear: does adding coprime Clifford + inter-qudit gates interpolate between classically simulable and universal circuits in a meaningful way?
-Current architectures typically use homogeneous registers—arrays of identical qubits or qudits—where the Clifford group plus inter-qudit entangling gates remains finite and classically simulable via stabilizer methods [Gottesman1998, AaronsonGottesman2004, Campbell2012], so universality requires magic state distillation [Bravyi2005].
-We propose an alternative: heterogeneous coprime architectures, where subsystems of dimensions $d_{1}$ and $d_{2}$ with $\gcd(d_{1},d_{2})=1$ achieve universality using only local Clifford gates and native inter-qudit, coupling.
-This principle could be realized in trapped ions [RingbauerEtAl2022, Low2020], superconducting circuits [Chicco_2023, Morvan2021], and photonic systems [Wang2020review, liu2024].
-The main advantage is eliminating magic state distillation overhead, which dominates resource costs in standard fault-tolerant schemes [Litinski2019, Gidney_2021], even with recent improvements [PrakashSaha2025].
-A potential drawback is increased control complexity for heterogeneous registers, but similar challenges already arise in modular and hybrid architectures.
-Finally, we note that the trichotomy discovered here may be the first instance of a broader phenomenon.
-Whenever a mathematical structure admits a notion of “dimension” with nontrivial factorization properties, similar classifications may arise.
-Exploring such analogies-in continuous-variable systems, in topological quantum computation, or in other algebraic settings-offers an interesting direction for future work.
+The results of this paper establish a classification of single-qudit universality, when supplementing the Clifford group, based on the prime factorization of the local Hilbert space dimension $d$.
+Depending on the value of $d$, a trichotomy arises capturing a hierarchy of algebraic obstructions within the Clifford group: maximality and irreducibility for prime dimensions, reducibility in the case of prime-power dimensions, and automatic resolution via coprime coupling in composite dimensions.
+Let us summarize the main findings.
+The Clifford group $\mathcal{C}_{d}$ is a maximal finite subgroup of $\mathrm{PSU}(p)$ and has irreducible adjoint representation if and only if $d$ is a prime (Theorem 25 and Theorem 24).
+This maximality and irreducibility imply that any non-Clifford gate—whether diagonal or permutation—suffices for universality (Corollary 27).
+For prime-power $d=p^{m}$ with $m\geq 2$, the adjoint representation becomes reducible (Theorem 30), fragmenting the Lie algebra into invariant subspaces indexed by divisibility.
+Universality requires gates that restore irreducibility: either diagonal gates $T_{s}$ with $s\nmid d$ satisfying the bound of Theorem 33, or non-affine permutations such as the transposition (0 1)(0\;1) analyzed in Theorem 39.
+For coprime composite $d=d_{1}\cdots d_{k}$, the intra-qudit CNOTs connecting subsystems of coprime dimension automatically generate the required non-Clifford resources via Bézout’s identity (Proposition 47).
+We emphasize that non-affine permutations constitute a distinct class of universal resources, available only for $d\geq 4$.
+For $d\in\{2,3\}$, every permutation is affine and, hence, a Clifford gate (Remark 10).
+This explains why diagonal magic gates remain necessary in qubit and qutrit systems, while for $d\geq 4$ universality can be achieved using only Clifford gates and permutations.
+The present work focuses on single-qudit universality, where the Clifford group is associated to the abelian group $\mathbb{Z}_{d}$.
+From single-qudit universality, general quantum universality follows via a result of Brylinski and Brylinski (Theorem 2).
+A natural extension concerns multi-qudit Clifford groups associated to $\mathbb{Z}_{d}^{n}$, corresponding to $n$ qudits of dimension $d$.
+For $d$ prime, the maximality of the $n$-qudit Clifford group was established by Nebe, Rains, and Sloane [34] using $\ell$-adic methods and the classification of finite simple groups.
+It would be interesting to determine whether the trichotomy identified here extends to the multi-qudit setting for general $d$, and whether the structure of $\mathbb{Z}_{d}^{n}$ for composite $d$ introduces new phenomena beyond those captured by the single-qudit analysis presented here.
+From a resource-theoretic perspective, existing measures of magic [2, 38, 22, 6] were developed primarily for qubit systems with explicit $T$-gate injection.
+The fact that permutation gates—which are classical reversible operations—serve as universal resources for $d\geq 4$ suggests that the relationship between magic and non-Cliffordness may be more subtle in higher-dimensional systems.
+On the experimental side, platforms accessing non-binary levels—trapped ions [35, 30], superconducting circuits [10, 33], and photonic systems [41, 28]—provide natural settings for implementing higher-dimensional quantum computation.
+Our results show that for $d\geq 4$, universality can be achieved without resorting to explicit diagonal magic injection: either via non-affine permutations (for any $d\geq 4$) or through intra-qudit CNOTs (for coprime composite $d$).
+A direction for future work is to compare the overhead associated with these high‑dimensional mechanisms against that of traditional magic‑state‑based approaches in low‑dimensional systems, such as magic state distillation [27, 16] or the recently proposed magic state cultivation [17].
+Determining whether these alternative routes reduce fault‑tolerant resource costs ultimately hinges on the relative difficulty of implementing and controlling such gates.
+We leave such an analysis for future research.
 Acknowledgements.
+The authors thank Vsevolod I.
+Yashin and Xingshan Cui for their valuable comments and suggestions.
 C.G.